### v1

# Two-Edge Connectivity via Pac-Man Gluing

Abstract
We study the 2-edge-connected spanning subgraph (2-ECSS) problem: Given a graph $G$, compute a connected subgraph $H$ of $G$ with the minimum number of edges such that $H$ is spanning, i.e., $V(H)=V(G)$, and $H$ is 2-edge-connected, i.e., $H$ remains connected upon the deletion of any single edge, if such an $H$ exists. The $2$-ECSS problem is known to be NP-hard. In this work, we provide a polynomial-time $(\frac{5}{4}+\varepsilon)$-approximation for the problem for an arbitrarily small $\varepsilon>0$, improving the previous best approximation ratio of $\frac{13}{10}+\varepsilon$.
Our improvement is based on two main innovations: First, we reduce solving the problem on general graphs to solving it on structured graphs with high vertex connectivity. This high vertex connectivity ensures the existence of a 4-matching across any bipartition of the vertex set with at least 10 vertices in each part. Second, we exploit this property in a later gluing step, where isolated 2-edge-connected components need to be merged without adding too many edges. Using the 4-matching property, we can repeatedly glue a huge component (containing at least 10 vertices) to other components. This step is reminiscent of the Pac-Man game, where a Pac-Man (a huge component) consumes all the dots (other components) as it moves through a maze. These two innovations lead to a significantly simpler algorithm and analysis for the gluing step compared to the previous best approximation algorithm, which required a long and tedious case analysis.
1 Introduction
We study the 2-edge-connected spanning subgraph (2-ECSS) problem, which is a fundamental fault-tolerant network design problem. In the 2-ECSS problem, we are given a graph $G$, and the objective is to compute a connected subgraph $H$ of $G$ with the minimum number of edges such that $H$ is spanning (i.e., $V(H)=V(G)$) and $H$ is 2-edge-connected (i.e., $H$ remains connected upon the deletion of any single edge: for all edges $e\in E(H)$, $H\setminus\{e\}$ is connected), if such an $H$ exists. While computing a (1-edge) connected spanning subgraph with the minimum number of edges—a spanning tree—admits simple polynomial-time algorithms, the 2-ECSS problem is MAX-SNP hard [13, 15]. In particular, assuming $\mathrm{P}\neq\mathrm{NP}$, there is no PTAS for the 2-ECSS problem. Consequently, there has been a quest to obtain good polynomial-time approximations for the problem.
A $2$-approximation for the $2$-ECSS problem is easily achieved by augmenting a DFS tree by picking the highest back edges for each non-root vertex. The first non-trivial approximation achieving a ratio of $\nicefrac{{3}}{{2}}$ was provided by Khuller and Vishkin [26]. Subsequently, the approximation ratio was first improved to $\nicefrac{{17}}{{12}}$ by Cheriyan, Sebö, and Szigeti [10], and later to $\nicefrac{{4}}{{3}}$ independently by Hunkenschröder, Vempala, and Vetta [37, 24] and Sebö and Vygen [34] using very different techniques. Several failed attempts (see Section F of [18]) were made to improve the approximation ratio of $\nicefrac{{4}}{{3}}$, making it a natural approximation barrier. In a breakthrough, Garg, Grandoni, and Jabal Ameli [18] obtained an approximation ratio of $\nicefrac{{118}}{{89}}+\varepsilon<1.326$. Subsequently, Kobayashi and Noguchi [27] observed that if a minimum $2$-edge cover that does not contain any isolated triangles can be computed in polynomial-time, then in fact the analysis of [18] will result in an approximation ratio of $1.3+\varepsilon$. They were able to show that such a triangle-free $2$-edge cover is polynomial-time computable using a polynomial-time computable maximum triangle-free 2-matching, a result proved in Hartvigsen’s PhD thesis [22], which was recently published in a journal [23] and subsequently improved and simplified by Paluch [32]. Additionally, a PTAS for triangle-free $2$-matching is available [4], which also results in a $(1.3+\varepsilon)$-approximation for the 2-ECSS problem.
1.1 Our Result
In this work, we improve the approximation ratio for the 2-ECSS problem to $\nicefrac{{5}}{{4}}+\varepsilon$.
Theorem 1.
For all $\varepsilon>0$, there exists a deterministic polynomial-time approximation algorithm for the $2$-ECSS problem with an approximation ratio of $\frac{5}{4}+\varepsilon$.
The previous best approximation ratio of $1.3+\varepsilon$ [27] depends on [18], a long and a highly involved paper consisting of tedious case analysis, making the result difficult to verify. In this work, we address this shortcoming by providing a simpler proof of a better approximation ratio of $\nicefrac{{5}}{{4}}+\varepsilon$.
1.2 Our Technique
Our proof is broadly based on the well-established template for such problems (as used in some form or another in [6, 7, 37, 24, 14, 29, 3]). It consists of four main steps: reducing the input instance to a structured instance, computing in polynomial time an object $\mathcal{O}$ (a minimum 2-edge cover with no isolated triangles) of the structured instance that is closely related to and costs less than an optimal desired object (minimum 2-ECSS), then converting $\mathcal{O}$ into a feasible solution (2-ECSS) by first covering the bridges in each connected component, followed by gluing the various connected components together while keeping the additional cost of the conversion to a small factor $c$ of the cost of $\mathcal{O}$. This results in a $(1+c)$-approximation. We will elaborate on these steps in detail in the following sections.
In brief, for the first three steps, we closely follow [18], which is an uncomplicated part of their paper, along with [27] for the second step, which helps in eliminating isolated triangles, an important fact without which the gluing would be cumbersome. Our main insight in this part is to make the input even more structured than in [18] by getting rid of all large 3-vertex cuts. This allows us to obtain a 4-matching lemma, which states that for any vertex partition of a structured input into two parts consisting of at least 10 vertices each, there exists a matching of size 4 going between the two parts. This lemma is then exploited in the gluing step. In the gluing step, we first glue a few components to obtain a huge component (consisting of at least 10 vertices). Then, with the aid of the 4-matching lemma, we are repeatedly able to glue a huge component with the other components, while keeping $c$ to be $\nicefrac{{1}}{{4}}$. This step is reminiscent of the popular Pac-Man game, where a Pac-Man (a huge component) eats up all the dots (other components) as it moves through a maze (component graph). Like in [18], we lose an additional $\varepsilon$ approximation factor in the reduction step itself, giving us an approximation ratio of $1+\nicefrac{{1}}{{4}}+\varepsilon=\nicefrac{{5}}{{4}}+\varepsilon$.
1.3 Additional Related Work
The 2-ECSS problem is closely related to augmentation problems, where we are given a graph having a specific structure (e.g., spanning tree, matching, forest, cactus, or disjoint paths) along with a set of additional edges called links. The task is to add the minimum number of links to the structure so that the resulting graph becomes spanning and 2-edge connected. For the (spanning) tree augmentation problem, several better than $2$-approximations are known [1, 5, 8, 9, 12, 14, 16, 21, 28, 29, 30, 31, 35, 36], with 1.393 as the currently best known approximation ratio by Cecchetto et al. [5]. For the forest augmentation problem, the first non-trivial approximation was obtained by Grandoni et al. [20] with a ratio of 1.9973. For the matching augmentation problem the current best approximation ratio is $\nicefrac{{13}}{{8}}$ [19] by Garg et al., while other non-trivial approximations are also known [2, 6, 7]. All the above problems have a natural weighted version. For the weighted 2-ECSS problem, obtaining an approximation ratio better than 2 is a major open problem. For the weighted tree augmentation problem the current best approximation ratio is $1.5+\varepsilon$ by Zenklusen and Traub [36]. The k-ECSS problem for $k\geq 2$ is another natural generalization of the 2-ECSS problem (see for e.g., [11, 17]).
1.4 Organization of the Paper
The paper proceeds as follows. In Section 2 we start with some preliminaries consisting of some basic definitions and notations. Subsequently, we tackle the four main steps of our approach in the following four sections: reduction to structured graphs in Section 3, obtaining a canonical 2-edge cover in Section 4, bridge covering in Section 5, and finally, we present the gluing step, the crux of this work, in Section 6. Certain proofs have been moved to the Appendix.
2 Preliminaries
We use standard graph theory notation. For a graph $G$, $V(G)$ and $E(G)$ refers to its vertex and edge sets, respectively. We use $\deg_{G}(v)$ to represent the degree of a vertex $v$ in the graph $G$ (when $G$ is clear from the context, we sometimes omit the subscript $G$). A pendant vertex in a graph is a vertex that is adjacent to exactly one other vertex. For a graph $G$, and a vertex set $S\subseteq V(G)$, $G[S]$ refers to the subgraph of $G$ induced on the vertex set $S$. We use components of a graph to refer to its maximal connected components. A cut vertex in a graph refers to a vertex in the graph, so that deleting it results in increasing the number of components of the graph. A $k$-vertex cut in a graph for $k\geq 1$ refers to a set of $k$ vertices such that deleting them results in increasing the number of components of the graph. A graph with no cut vertex is called 2-vertex connected (2VC). Similarly, a bridge in a graph refers to an edge in the graph, so that deleting it results in increasing the number of components in the graph. A connected graph without bridges is called 2-edge connected (2EC). We say a subgraph $H$ of a graph $G$ is spanning if $V(H)=V(G)$.
A spanning subgraph $H$ of a graph $G$ is called a 2-edge cover of $G$ if the degree of each vertex in $H$ is at least 2. A cycle refers to a simple cycle. We use $\mathcal{C}_{i}$ to represent a cycle on $i$ vertices and refer to $\mathcal{C}_{3}$ as a triangle. We say a 2-edge cover is triangle-free, if none of its components is a triangle.
Given a graph $G$ and a vertex set $S\subseteq V(G)$, $G|S$ denotes the contracted graph obtained from contracting the vertex set $S$ into a single vertex. Graph contraction may give rise to parallel edges and self-loops. The edges of $G$ and $G|S$ are in one-to-one correspondence. For a graph $G$ and a subgraph $H$, we use $G|H$ to denote $G|V(H)$.
For a graph $G$, we use $|G|$ to denote $|E(G)|$. For a 2EC graph $G$ we use $\mathsf{OPT}(G)$ to denote a 2-ECSS of $G$ with the minimum number of edges. We define $\mathsf{opt}(G)\coloneq|\mathsf{OPT}(G)|$ (when $G$ is clear from the context we may just say $\mathsf{opt}$ and $\mathsf{OPT}$ instead). For a graph $G$, a subgraph $H$ is called a 2-ECSS of $G$ if it is 2EC and spanning.
3 Reduction to Structured Graphs
In our approximation algorithm for the 2-ECSS problem we will assume that the input is 2EC, otherwise it has no feasible solutions, which is easy to check: delete each edge and check if the graph is connected. Note that the input graph may contain parallel edges or self loops. In [18], structured graphs were defined, and it was shown that, for each small enough $\varepsilon>0$, if the 2-ECSS problem can be approximated in polynomial time on structured graphs with an approximation ratio $\alpha\geq\nicefrac{{6}}{{5}}$, then it can be approximated in polynomial time on general graphs with only a $2\varepsilon$ loss in approximation, i.e., with an approximation ratio of $\alpha+2\varepsilon$. In this work, we define structured graphs to be even more restricted than in [18], and prove a similar approximation-preserving reduction up to an $O(\varepsilon)$ loss.
Informally, a structured graph is a simple graph that is 2VC with sufficiently many vertices and does not contain any contractible subgraphs, irrelevant edges, non-isolating 2-vertex cuts, or large 3-vertex cuts. Forbidding large 3-vertex cuts in the definition is the only difference compared to the definition in [18]. We now define the various terms appearing in the definition of structured graphs before defining structured graphs formally.
Definition 1 ($\alpha$-contractible subgraph [18]).
Let $\alpha\geq 1$ be a fixed constant. A 2EC subgraph $C$ of a 2EC graph $G$ is $\alpha$-contractible if every 2-ECSS of $G$ contains at least $\frac{1}{\alpha}\lvert E(C)\rvert$ edges with both endpoints in $V(C)$.
The intuition behind defining $\alpha$-contractible subgraphs is the following. While designing an $\alpha$-approximation algorithm for the 2-ECSS problem, if we can find an $\alpha$-contractible subgraph $C$ in polynomial-time, we can contract $C$ into a single vertex and find an $\alpha$-approximate solution on $G|C$, and add the edges of $C$ to get an $\alpha$-approximate solution for $G$. The approximation guarantee holds since $\mathsf{opt}(G)\geq\mathsf{opt}(G|C)+\mathsf{opt}(C)$, which is true since $\mathsf{OPT}(G)$ needs to have at least $\nicefrac{{1}}{{\alpha}}\cdot\lvert E(C)\rvert$ from $G[V(C)]$. Thus, in our approximation-preserving reduction all constant-sized $\alpha$-contractible subgraphs will be identified and handled.
Definition 2 (irrelevant edge [18]).
Given a graph $G$ and an edge $e=uv\in E(G)$, we say that $e$ is irrelevant if $\{u,v\}$ is a 2-vertex cut of $G$.
Definition 3 (non-isolating 2-vertex cut [18]).
Given a graph $G$, and a 2-vertex cut $\{u,v\}\subseteq V(G)$ of $G$, we say that $\{u,v\}$ is isolating if $G\setminus\{u,v\}$ has exactly two connected components, one of which is of size 1. Otherwise, it is non-isolating.
Note that getting rid of non-isolating cuts means that $G$ can be assumed to be almost 3-vertex-connected. In this work, we introduce the definitions of small and large 3-vertex cuts.
Definition 4 (small and large 3-vertex cuts).
Given a graph $G$ and a 3-vertex cut $\{u,v,w\}\subseteq V(G)$ of $G$, we say that $\{u,v,w\}$ is small if $G\setminus\{u,v,w\}$ has exactly two connected components such that one of them has at most 6 vertices. Otherwise, we call it large.
We are now ready to define structured graphs formally.
Definition 5 (($\alpha,\varepsilon)$-structured graph).
Given $\alpha\geq 1$ and $\varepsilon>0$, a graph $G$ is $(\alpha,\varepsilon)$-structured if it is simple, 2VC, it contains at least $\frac{4}{\varepsilon}$ vertices, and it does not contain
  1. 1.
    $\alpha$-contractible subgraphs of size at most $\frac{2}{\varepsilon}$,
  2. 2.
    irrelevant edges,
  3. 3.
    non-isolating 2-vertex cuts, and
  4. 4.
    large 3-vertex cuts.
In [18], it was shown how to get rid of parallel edges, self loops, irrelevant edges, cut-vertices, small $\alpha$-contractible subgraphs and non-isolating 2-vertex cuts via an approximation-preserving reduction. We add to their proof and show that we can also get rid of large 3-vertex cuts by generalizing their proof for non-isolating 2-vertex cuts in various ways. This is a significant part of this work. In Appendix A we prove the following approximation-preserving reduction. This proof is a bit long, but it admits a systematic case analysis. After understanding the main ideas through the first few cases, the remaining cases can actually be left as a simple exercise (though for completeness, we cover all the cases). {restatable}lemmalemmaReduction For all $\alpha\geq\frac{5}{4}$ and $\varepsilon\in(0,\frac{1}{24}]$, if there exists a deterministic polynomial-time $\alpha$-approximation algorithm for $2$-ECSS on ($\alpha,\varepsilon)$-structured graphs, then there exists a deterministic polynomial-time $(\alpha+4\varepsilon)$-approximation algorithm for $2$-ECSS.
When $\alpha$ and $\varepsilon$ are clear from the context, we simply write contractible and structured instead of $\alpha$-contractible and $(\alpha,\varepsilon)$-structured, respectively.
Benefits of stuctured graphs. After we have the above reduction, we need to design approximation algorithms only for structured graphs. How do we exploit the fact that the input graph is structured? We will make generous use of the fact that there are no small contractible subgraphs. Furthermore, it was shown in [18] that certain 3-matchings exist in a structured graph, which will be very useful.
Lemma 2 ($3$-Matching Lemma [18]).
Let $G=(V,E)$ be a $2$VC simple graph without irrelevant edges and without non-isolating 2-vertex cuts. Consider any partition $(V_{1},V_{2})$ of $V$ such that for each $i\in\{1,2\}$, $|V_{i}|\geq 3$, and if $|V_{i}|=3$, then $G[V_{i}]$ is a triangle. Then, there exists a matching of size $3$ between $V_{1}$ and $V_{2}$.
We further make use of the non-existence of large $3$-vertex cuts in structured graphs to establish a $4$-matching-lemma, which is going to be very useful in the gluing step. [^1] In fact, with similar techniques we can also get rid of large $k$-vertex cuts (for any constant $k$), where both sides of the partition have $\Omega(\frac{k}{1-\alpha})$ many vertices, leading to a similar $k$-Matching Lemma. However, we do not see an easy way to exploit this fact in improving the approximation ratio with the current techniques.
Lemma 3 (4-Matching Lemma).
Let $G$ be a graph without large 3-vertex cuts. Consider any partition $(V_{1},V_{2})$ of $V(G)$ such that for each $i\in\{1,2\}$, $\lvert V_{i}\rvert\geq 10$. Then, there exists a matching of size $4$ between $V_{1}$ and $V_{2}$ in $G$.
Proof.
Consider the bipartite graph $F$ induced by the edges with exactly one endpoint in $V_{1}$ (and the other in $V_{2}$). For contradiction, we assume that a maximum matching $M$ in $F$ has size $\lvert M\rvert\leq 3$. Then, by the König Egerváry Theorem (see, e.g., [33]), there exists a vertex cover $C$ of $F$ of size $\lvert M\rvert$. Since $C$ is a vertex cover of $F$, there are no edges in $F$ (hence in $G$) between $U_{1}\coloneqq V_{1}\setminus C$ and $U_{2}\coloneqq V_{2}\setminus C$. Therefore, $C$ is a 3-vertex cut in $G$, which is large as $\lvert U_{1}\rvert,\lvert U_{2}\rvert\geq 7$ (because $\lvert V_{1}\rvert,\lvert V_{2}\rvert\geq 10$ and $\lvert C\rvert\leq 3$), a contradiction. ∎
4 Canonical 2-Edge Cover
It was shown in [27] that a minimum triangle-free 2-edge cover can be computed in polynomial time.
Lemma 4 (Proposition 7 in [27]).
For a graph $G=(V,E)$, we can compute a triangle-free 2-edge cover of $G$ with minimum cardinality in polynomial time if one exists.
Notice that for our input structured graph $G$, a 2-ECSS of $G$ is also a triangle-free 2-edge cover. Thus, if $H$ is a minimum triangle-free 2-edge cover of our input $G$, then $\lvert H\rvert\leq\mathsf{opt}(G)$. Roughly speaking, we will compute $H$ using the above lemma, and then transform $H$ into a 2-ECSS of $G$ by adding at most $\nicefrac{{1}}{{4}}\cdot\lvert H\rvert$ edges to it. Thus, we would get a 2-ECSS with at most $\nicefrac{{5}}{{4}}\cdot\mathsf{opt}(G)$, proving our main result. In reality, we might add some more edges and delete some edges, so that the net total of additional edges will be at most $\nicefrac{{1}}{{4}}\cdot\lvert H\rvert$.
We will first convert $H$ into a canonical triangle-free 2-edge cover without changing the number of edges in it. In order to define canonical we first need the following definitions pertaining to a 2-edge cover.
Definition 6.
Let $H$ be a $2$-edge cover of some graph. $H$ is bridgeless if it contains no bridges, i.e., all the components of $H$ are $2$-edge connected. A component of $H$ is complex if it contains a bridge. Maximal222w.r.t. the subgraph relationship $2$-edge-connected subgraphs of $H$ are blocks. A block $B$ of some complex component $C$ of $H$ is called pendant if $C\setminus V(B)$ is connected. Otherwise it is non-pendant. Note that by contracting each block of a complex component to a single vertex, we obtain a tree whose leaves correspond to pendant blocks.
Definition 7 (canonical).
A $2$-edge cover $H$ is canonical if the following properties are satisfied:
  • Each non-complex component of $H$ is either a $\mathcal{C}_{i}$, for $4\leq i\leq 7$, or contains at least $8$ edges.
  • For every complex component, each of its pendant blocks contains at least $6$ edges and each of its non-pendant blocks contains at least $4$ edges.
We establish the following lemma in Appendix B that shows how to convert a minimum triangle-free 2-edge cover of a structured graph into a minimum canonical 2-edge cover in polynomial time.
{restatable}
lemmalemmaCanonicalMain Given a structured graph $G$, in polynomial time we can compute a canonical $2$-edge cover $H$ of $G$, where the number of edges in $H$ is at most (and hence equal to) the number of edges in a minimum triangle-free $2$-edge cover. This definition of canonical and the proof of the above lemma are similar to the ones in [18], except that now components with $7$ edges have to be a $\mathcal{C}_{7}$.
5 Bridge Covering and Credits
We now have a structured graph $G$ and a minimum canonical 2-edge cover $H$ of $G$ such that $|H|\leq\mathsf{opt}(G)$. We will show how to convert $H$ into a 2-ECSS of $G$ by adding a net total of at most $\nicefrac{{1}}{{4}}\cdot\lvert H\rvert$ edges to it. There are two reasons why $H$ is not already a 2-ECSS of $G$: First, $H$ might have multiple components. Second, some of these components might have bridges. We will first modify $H$ so that the resulting components do not have any bridges. Then, we will glue the various resulting 2EC components together to get a desired 2-ECSS, which we defer to the next section. Throughout this transformation, we will keep the property that the intermediate solution is a canonical 2-edge cover.
To keep track of how many edges we are adding, we will define a credit invariant that is satisfied throughout the transformation and a cost function. We have a total budget of $\nicefrac{{1}}{{4}}\cdot\lvert H\rvert$, which we will distribute all over $H$ as credits of different kinds. The cost of $H$ is defined keeping in mind that the cost of a single edge is 1. $\mathrm{cost}(H)$ is the total number of edges in $H$ plus the total credits in $H$. Thus, initially the $\mathrm{cost}(H)$ will be at most $\nicefrac{{5}}{{4}}\cdot\mathsf{opt}$. At each step of our transformation, we aim to ensure that the cost never increases, so that the final solution we obtain will have cost at most $\nicefrac{{5}}{{4}}\cdot\mathsf{opt}$, and we will be done. We now define our credit invariant and the cost function, which is also used in the next section.
Definition 8 (credits).
Let $H$ be a $2$-edge cover of $G$. We keep the following credit invariant for blocks, components and bridges:
  • Each 2EC component $C$ of $H$ that is a $\mathcal{C}_{i}$, for $4\leq i\leq 7$, receives credit $\mathrm{cr}(C)=\frac{1}{4}|E(C)|$.
  • Each 2EC component $C$ of $H$ that contains 8 or more edges receives credit $\mathrm{cr}(C)=2$.
  • Each block $B$ of a complex component $C$ of $H$ receives a credit $\mathrm{cr}(B)=1$.
  • Each bridge $e$ of a complex component $C$ of $H$ receives credit $\mathrm{cr}(e)=\frac{1}{4}$.
  • Each complex component $C$ of $H$ receives a component credit $\mathrm{cr}(C)=1$.
Definition 9 (cost).
The cost of a $2$-edge cover $H$ of $G$ is defined as $\mathrm{cost}(H)=|H|+\mathrm{cr}(H)$. Here $\mathrm{cr}(H)$ denotes the sum total of all credits in $H$.
We first show that we can distribute our budget of $\nicefrac{{1}}{{4}}\cdot\lvert H\rvert$ over our minimum canonical triangle-free 2-edge cover to satisfy the credit invariant. In other words, we assume the credits are there as required by the credit invariant and show that $\mathrm{cost}(H)\leq\nicefrac{{5}}{{4}}\cdot\lvert H\rvert$.
Lemma 5.
Let $H$ be a canonical $2$-edge cover of a graph $G$. Then, we have $\mathrm{cost}(H)\leq\frac{5}{4}|H|$.
Proof.
We show that $\mathrm{cost}(C)\leq\frac{5}{4}|C|$ for each component $C$ in $H$. This then implies that $\mathrm{cost}(H)\leq\frac{5}{4}|H|$.
First, consider a 2EC component $C$ that is a $\mathcal{C}_{i}$, for $4\leq i\leq 7$. Since $C$ has exactly $i$ edges, $\mathrm{cost}(C)=|C|+\frac{1}{4}|E(C)|=\frac{5}{4}|C|$. Second, consider a 2EC component $C$ with at least 8 edges. Note that $\frac{1}{4}|E(C)|\geq 2$. Hence, we have that $\mathrm{cost}(C)=|C|+2\leq|C|+\frac{1}{4}|E(C)|=\frac{5}{4}|C|$.
Third, consider a complex component $C$. Each bridge $e$ of the complex component receives credit $\mathrm{cr}(e)=\frac{1}{4}$. Each block receives a credit of $1$. Finally, the component credit of $C$ is 1. Note that since $H$ is canonical, each pendant block has at least 6 edges, and non-pendant blocks have at least 4 edges. Moreover, there are at least 2 pendant blocks in $C$. Let $c_{p}$ be the number of pendant blocks of $C$, $c_{n}$ be the number of non-pendant blocks of $C$, and $c_{b}$ be the number of bridges in $C$.
Note that $\lvert C\rvert\geq c_{b}+6c_{p}+4c_{n}$. Therefore, $\frac{1}{4}\lvert C\rvert\geq\frac{1}{4}c_{b}+\frac{3}{2}c_{p}+c_{n}$. Since $c_{p}\geq 2$ implies $\frac{3}{2}c_{p}\geq c_{p}+1$, we have $\frac{1}{4}\lvert C\rvert\geq\frac{1}{4}c_{b}+c_{p}+c_{n}+1$. Now, we plug this fact in the cost calculation:
$\mathrm{cost}(C)=|C|+\mathrm{cr}(C)=|C|+\frac{1}{4}c_{b}+c_{p}+c_{n}+1\leq|C|+\frac{1}{4}|C|\leq\frac{5}{4}|C|.$
We are now ready to state our main lemma for bridge covering, which essentially is already proved in [18]. We add some comments in Appendix C on how to adapt the proof to our setting.
{restatable}
lemmalemmaBridgeCoveringMain There is a polynomial-time algorithm that takes as input a canonical 2-edge cover $H$ of a structured graph $G$ and outputs a bridgeless canonical 2-edge cover $H^{\prime}$ of $G$ such that $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)$.
We will feed our minimum canonical 2-edge cover $H$ to the algorithm in the above lemma to obtain a bridgeless canonical 2-edge cover $H^{\prime}$ and $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)\leq\nicefrac{{5}}{{4}}\cdot\mathsf{opt}(G)$. Each component of $H^{\prime}$ now is 2EC and either has $8$ or more edges, in which case it has a credit of $2$, otherwise it is a cycle with $i$ edges ($\mathcal{C}_{i}$) for $4\leq i\leq 7$ with $\nicefrac{{1}}{{4}}\cdot i$ credits. We will use these credits to glue the 2EC components of $H^{\prime}$ together to obtain a desired 2-ECSS of $G$, as explained in the next section.
6 Gluing
We have a structured graph $G$ as input and a bridgeless canonical 2-edge cover $H$ of $G$ with $\mathrm{cost}(H)\leq\nicefrac{{5}}{{4}}\cdot\mathsf{opt}(G)$ such that $H$ satisfies the credit invariant: $2$ in components with at least $8$ edges, and $\nicefrac{{i}}{{4}}$ in components that are cycles with $i$ edges, for $4\leq i\leq 7$. Now, in the final step of our algorithm, we glue the components of $H$ together (while maintaining the credit-invariant) using the credits available and a marginal additional cost of $3$. Specifically, we show the following.
Lemma 6.
Given a structured graph $G$ and a bridgeless canonical 2-edge cover $H$ of $G$, we can transform $H$ in polynomial time into a 2-ECSS $H^{\prime}$ of $G$ such that $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)+3$.
Using the above lemma, the proof of Theorem 1 follows easily: For a structured graph $G$, we output the 2-ECSS $H^{\prime}$ of $G$ obtained by applying the lemma above. We know $\lvert H^{\prime}\rvert+\mathrm{cr}(H^{\prime})=\mathrm{cost}(H^{\prime})\leq\frac{5}{4}\cdot\mathsf{opt}(G)+3$. Now $\mathrm{cr}(H^{\prime})=2$, as it will be a single 2EC component with more than $8$ edges. Thus, $\lvert H^{\prime}\rvert\leq\frac{5}{4}\cdot\mathsf{opt}(G)+1=(\frac{5}{4}+\frac{\varepsilon}{4})\cdot\mathsf{opt}(G)$, where the last inequality follows from $\mathsf{opt}(G)\geq\lvert V(G)\rvert\geq\frac{4}{\varepsilon}$, which holds for structured graphs $G$. Now, using the approximation preserving reduction, Definition 5, for general graphs $G$, we incur an additional loss of $\varepsilon$ to get an approximation ratio of $\frac{5}{4}+\frac{5}{4}\varepsilon=\frac{5}{4}+\varepsilon^{\prime}$, where $\varepsilon^{\prime}=\frac{5}{4}\varepsilon$, proving Theorem 1.
It remains to prove Lemma 6. To this end, we first introduce some definitions. In what follows, unless otherwise mentioned, $G$ is a structured graph and $H$ is a bridgeless canonical 2-edge cover of $G$.
Definition 10 (large and huge).
We say that a component $C$ of $H$ is large if $|V(C)|\geq 8$ and we say that it is huge if $|V(C)|\geq 10$.
Notice that huge components are also large and must have 2 credits. The benefit of having a huge component is that the 4-matching lemma can be applied to it to conclude that each huge component has a matching of size $4$ going out of it (whenever there are at least 10 vertices outside the huge component).
Definition 11 (component graph).
The component graph ^$\bm{\hat{}}{G}_{H}$ w.r.t $G$ and $H$ is the multigraph obtained from $G$ by contracting the vertices of each 2EC component of $H$ into a single node. We call the vertices of ^$\bm{\hat{}}{G}_{H}$ nodes to distinguish them from the vertices in $G$. There is a one-to-one correspondence between the edges of ^$\bm{\hat{}}{G}_{H}$ and $G$. Furthermore, there is a one-to-one correspondence between the nodes of ^$\bm{\hat{}}{G}_{H}$ and the components of $H$. We will denote the component corresponding to a node $A$ by $C_{A}$.
Although not strictly required, to remove any ambiguities in our presentation later while dealing with cycles, we get rid of all the self-loops that appear in the component graph, and call the resulting graph the component graph. Now, observe that the component graph $\hat{G}_{H}$ is 2EC as $G$ is 2EC (as contracting a set of vertices in a 2EC graph results in a 2EC graph; also deleting self loops from a 2EC graph results in a 2EC graph).
Definition 12 (non-trivial and trivial segments).
A non-trivial segment $S$ of ^$\bm{\hat{}}{G}_{H}$ is a maximal 2-node-connected subgraph of ^$\bm{\hat{}}{G}_{H}$ consisting of at least 3 nodes. A trivial segment is a single node of ^$\bm{\hat{}}{G}_{H}$ that is not part of any non-trivial segment. For a segment $S$, $V(S)$ refers to the set of all the vertices in the various components corresponding to the nodes contained in $S$.
Notice that we can break the component graph into segments such that two segments can potentially overlap only at a cut node. This is easily achievable by identifying all the cut nodes of the component graph.
Furthermore, since a node $A$, that is a trivial segment, is not part of any cycles of length at least 3 in ^$\bm{\hat{}}{G}_{H}$ (otherwise $A$ would be part of a non-trivial segment), $A$ has to be either a cut node of ^$\bm{\hat{}}{G}_{H}$ or a pendant node (neighbor of exactly one other node) of ^$\bm{\hat{}}{G}_{H}$. In either case, we can apply the 3-matching lemma, Lemma 2, to argue that $V(C_{A})$ has a matching of size 3 going out to $V(C_{B})$, where node $B$ is a neighbor of node $A$ in ^$\bm{\hat{}}{G}_{H}$. The condition of the lemma is satisfied by partitioning $V(G)$ into two parts such that one part has vertices from $V(C_{A})$ and the other part has vertices from $V(C_{B})$ and the only edges going across the parts are between $V(C_{A})$ and $V(C_{B})$. We have the following propositions.
Proposition 7.
Let $A$ be a trivial segment of ^$\bm{\hat{}}{G}_{H}$. Then, $A$ is either a cut node of ^$\bm{\hat{}}{G}_{H}$ or it is a pendant node of ^$\bm{\hat{}}{G}_{H}$.
Proposition 8.
Let $A$ and $B$ be adjacent nodes in ^$\bm{\hat{}}{G}_{H}$ such that $A$ is a trivial segment. Then, there exists a matching of size $3$ going between $V(C_{A})$ and $V(C_{B})$ in $G$.
Having set up the requisite definitions, we now outline the steps involved in the gluing process. First, if $H$ does not contain a huge component, then we first glue a few components together, so that we obtain a canonical bridgeless 2-edge cover $H^{\prime}$ that has a huge component, while maintaining the credit invariant. We will show that such an $H^{\prime}$ can be obtained with $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)+3$. After this step, we assume $H$ has a huge component, an invariant that will be maintained at each iteration of our algorithm. Next, let $L$ be a huge component in $H$. We will distinguish two cases: either $L$ is a trivial segment of ^$\bm{\hat{}}{G}_{H}$, or it is part of a non-trivial segment of ^$\bm{\hat{}}{G}_{H}$. In either case we show that we can make progress, i.e., we can find in polynomial time a canonical bridgeless 2-edge cover $H^{\prime}$ that has fewer components than $H$ such that $H^{\prime}$ contains a huge component and $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)$. Exhaustively applying the above step will result in a single bridgeless 2EC component, a 2-ECSS of $G$. Thus, if we can establish the existence of these steps, Lemma 6 follows immediately. These steps are encapsulated in the following three lemmas.
Lemma 9.
If H does not contain a huge component, then we can compute in polynomial time a canonical bridgeless 2-edge cover $H^{\prime}$ such that $H^{\prime}$ contains a huge component and $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)+3$.
Lemma 10.
Let $C_{L}$ be a huge component such that $L$ is a trivial segment of ^$\bm{\hat{}}{G}_{H}$ and $H\neq C_{L}$. Then, in polynomial time we can compute a canonical bridgeless 2-edge cover $H^{\prime}$ of $G$ such that $H^{\prime}$ has fewer components than $H$, $H^{\prime}$ contains a huge component, and $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)$.
Lemma 11.
Let $C_{L}$ be a huge component such that $L$ is part of a non-trivial segment of ^$\bm{\hat{}}{G}_{H}$. Then, in polynomial time we can compute a canonical bridgeless 2-edge cover $H^{\prime}$ of $H$ such that $H^{\prime}$ has fewer components than $H$, $H^{\prime}$ contains a huge component, and $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)$.
As explained above, the above three lemmas immediately imply Lemma 6. All that remains is to prove the above three lemmas, which we do in the next three subsections.
6.1 Proof of Lemma 9
Proof.
W.l.o.g, $H$ has at least two components; if $H$ consists of a single component, it has to be huge as $\lvert V(H)\rvert=\lvert V(G)\rvert\geq\frac{4}{\varepsilon}\geq 10$. Let $A$ be a node in ^$\bm{\hat{}}{G}_{H}$. Let $K$ be some cycle in ^$\bm{\hat{}}{G}_{H}$ containing $A$ which must exist (and easily found) as ^$\bm{\hat{}}{G}_{H}$ is 2EC and has at least 2 nodes.
Let $H_{1}\coloneq H\cup K$ (i.e., add the edges of the cycle $K$ to $H$). Observe that $H_{1}$ is a canonical bridgeless 2-edge cover with fewer components than $H$, and the newly created component, call it $C_{B}$, containing the vertices of $C_{A}$ is large ($\geq 8$ vertices: as each component of $H$ contains at least $4$ vertices and $|K|\geq 2$). Furthermore, $|H_{1}|=|H|+|K|$ and $\mathrm{cr}(H_{1})\leq\mathrm{cr}(H)-|K|+2$, because every component in $K$ has a credit of at least $1$ in $H$ and the newly created large component receives a credit of $2$. If $H_{1}$ consists of only a single component, namely $C_{B}$, then $C_{B}$ must be huge and we are immediately done.
Otherwise, arguing as before, we can find another cycle $K^{\prime}$ through $B$ in ^$\bm{\hat{}}{G}_{H_{1}}$ and add $K^{\prime}$ to $H_{1}$ to obtain $H_{2}$, i.e., $H_{2}\coloneq H\cup K\cup K^{\prime}$. Observe that $H_{2}$ is again a canonical bridgeless 2-edge cover and the newly formed component (containing the vertices of $B$) has at least 12 vertices, and is huge. We have $|H_{2}|=|H|+|K|+|K^{\prime}|$ and $\mathrm{cr}(H_{2})\leq\mathrm{cr}(H_{1})-(|K^{\prime}|+1)+2$ (since every component in $H_{1}$ has a credit of 1, $\mathrm{cr}(B)=2$, and the newly formed component receives 2 credits). Plugging the upper bound on $\mathrm{cr}(H_{1})$ from above, we have $\mathrm{cr}(H_{2})\leq\mathrm{cr}(H)-|K|+2-(|K^{\prime}|+1)+2=\mathrm{cr}(H)-|K|-|K^{\prime}|+3$. Thus, $\mathrm{cost}(H_{2})=|H_{2}|+\mathrm{cr}(H_{2})\leq|H|+|K|+|K^{\prime}|+\mathrm{cr}(H)-|K|-|K^{\prime}|+3=|H|+\mathrm{cr}(H)+3=\mathrm{cost}(H)+3$, which completes the proof. ∎
6.2 Proof of Lemma 10
Proof.
Let $C_{L}$ be the huge component such that the node $L$ is a trivial segment of ^$\bm{\hat{}}{G}_{H}$ and $H\neq C_{L}$. Our goal is to construct a canonical bridgeless 2-edge-cover $H^{\prime}$ of $G$ with fewer components compared to $H$ such that $H^{\prime}$ contains a huge component and $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)$.
Since $H\neq C_{L}$, ^$\bm{\hat{}}{G}_{H}$ has at least 2 nodes, and as ^$\bm{\hat{}}{G}_{H}$ is 2EC, $L$ has a neighbor, say $A$. By Proposition 8, we know that there must be a 3-matching $M_{3}$ between $C_{L}$ and $C_{A}$. First, note that if $C_{A}$ is a large component, we simply add any two edges of $M_{3}$ to $H$ to obtain $H^{\prime}$, and it can be easily checked that $H^{\prime}$ meets our goal. Next assume that $C_{A}$ is a $\mathcal{C}_{4}$ or a $\mathcal{C}_{5}$. In either case, there must be two distinct edges in $M_{3}$, say $e_{1}$ and $e_{2}$, whose endpoints in $C_{A}$ are adjacent via an edge $e$ in $C_{A}$. We set $H^{\prime}\coloneq(H\setminus\{e\})\cup\{e_{1},e_{2}\}$ and observe that $H^{\prime}$ meets our goal. The only cases that remain are $C_{A}$ is a $\mathcal{C}_{6}$ or a $\mathcal{C}_{7}$. We consider the following exhaustive cases to complete the proof.
Case 1: There are two distinct edges in $M_{3}$, say $e_{1}$ and $e_{2}$, which are incident to $u$ and $v$ in $C_{A}$ and there is a Hamiltonian Path $P$ of $G[V(C_{A})]$ from $u$ to $v$. We set $H^{\prime}\coloneq(H\setminus E(C_{A}))\cup\{e_{1},e_{2}\}\cup P$ and, like before, it is easy to check that $H^{\prime}$ meets our goal.
Case 2: $A$ is not a pendant node in ^$\bm{\hat{}}{G}_{H}$ (i.e., it has other neighbors other than $A$). Then, Lemma 3 is applicable, implying a 4-matching $M_{4}$ between $C_{A}$ and $C_{L}$. This is because we can partition $V(G)$ in two parts to have $V(C_{L})$ in one part and $V(C_{A})\cup V(C_{B})$ in another part (see Proposition 7) such that the only edges crossing the parts are between $V(C_{L})$ and $V(C_{A})$, and $|V(C_{A})|+|V(C_{B})|\geq 6+4=10$ and $|V(C_{L})|\geq 10$, because $L$ is huge. Arguing as before, there must exist edges $e_{1}$ and $e_{2}$ whose endpoints in $C_{A}$ are adjacent via an edge $e$ in $C_{A}$. Therefore, we conclude that actually Case 1 holds as $C_{A}\setminus\{e\}$ forms a Hamiltonian path of $G[V(C_{A})]$.
Case 3: $A$ is pendant node of ^$\bm{\hat{}}{G}_{H}$ (i.e., it has exactly one neighbor in ^$\bm{\hat{}}{G}_{H})$. Using the fact that we are not in Case 1 or Case 2, we prove that $C_{A}$ is contractible, which contradicts the fact that $G$ is structured, and hence Case 3 is redundant.
To prove that $C_{A}$ is contractible, let the vertices of the cycle $C_{A}$ be labelled as $v_{1}-v_{2}-v_{3}-v_{4}-v_{5}-v_{6}-v_{1}$ if $C_{A}$ is a $\mathcal{C}_{6}$ or $v_{1}-v_{2}-v_{3}-v_{4}-v_{5}-v_{6}-v_{7}-v_{1}$ if $C_{A}$ is a $C_{7}$. Since we are not in Case 1 and there is the 3-matching $M_{3}$ between $C_{L}$ and $C_{A}$, $C_{L}$ is adjacent to exactly three mutually non-adjacent vertices of $C_{A}$ via $M_{3}$, say w.l.o.g. $\{v_{1},v_{3},v_{5}\}$ (up to symmetry this is the only possibility in both the $\mathcal{C}_{6}$ and $\mathcal{C}_{7}$ cases). We say an edge in $G[V(C_{A})]$ is a nice diagonal if both its endpoints are in $\{v_{2},v_{4},v_{6}\}$. Therefore, the set of nice diagonals are $\{v_{2}v_{4},v_{4}v_{6},v_{6}v_{1}\}$. If there exists a nice diagonal $e$ in $G[V(C_{A})]$, then observe that there are two distinct edges in $M_{3}$ that are incident to say $u$ and $v$ in $C_{A}$ and there is a Hamiltonian Path $P\subseteq\{e\}\cup E(C_{A})$ from $u$ to $v$ in $G[V(C_{A})]$. For example, if the diagonal $v_{2}v_{4}$ exists in $G$ and $C_{A}$ is $\mathcal{C}_{6}$, then the desired Hamiltonian path is $v_{3}-v_{2}-v_{4}-v_{5}-v_{6}-v_{1}$ (for $\mathcal{C}_{6}$, up to symmetry, this is the only case; the $\mathcal{C}_{7}$ case is left to the reader to verify, a simple exercise). Hence, we are in fact in Case 1. Otherwise, if none of these nice diagonals are in $G$, observe that $C_{A}$ is contractible: The three sets of edges of $G$, namely, edges incident on $v_{2}$, edges incident on $v_{4}$, and edges incident on $v_{6}$ are all inside $G[V(C_{A})]$ and mutually disjoint. As any 2-ECSS of $G$ must contain at least $2$ edges incident on each vertex, it has to contain at least $6$ edges from $G[V(C_{A})]$, and hence $C_{A}$ is contractible. ∎
6.3 Proof of Lemma 11
We will make use of the following lemma, which we will repeatedly apply to glue a huge component with all the $\mathcal{C}_{4}$ and $\mathcal{C}_{5}$ components (and potentially also larger components) in a non-trivial segment.
Lemma 12.
Let $C_{L}$ be a huge component and $C_{A}$ be a $\mathcal{C}_{4}$ or a $\mathcal{C}_{5}$ component of $H$ such that $L$ and $A$ are in the same non-trivial segment $S$ of ^$\bm{\hat{}}{G}_{H}$. Given such an $L$, $A$, and $S$, we can find in polynomial time a cycle $K$ in ^$\bm{\hat{}}{G}_{H}$ through the nodes $L$ and $A$, where the two edges of the cycle $K$ entering the node $A$ are incident on two vertices $u$ and $v$ of $C_{A}$, such that at least one of the following two conditions holds.
  1. (a)
    We can find a Hamiltonian Path from $u$ to $v$ in $G[V(C_{A})]$.
  2. (b)
    We can find a node $D$ of ^$\bm{\hat{}}{G}_{H}$ outside of the cycle $K$ and a set of edges $F\subseteq E(G)$ disjoint to $E(K)$ such that $|F|=|E(C_{A})|+|E(C_{D})|$ and $F\cup\{uv\}$ forms a 2-ECSS of $G[V(C_{A})\cup V(C_{D})]$.
Using the above lemma, we prove Lemma 11, and then prove Lemma 12.
Proof of Lemma 11.
We are given a non-trivial segment $S$ of ^$\bm{\hat{}}{G}_{H}$ which has a node $L$ such that $C_{L}$ is a huge component. Our goal is to construct a canonical bridgeless 2-edge-cover $H^{\prime}$ of $G$ with fewer components compared to $H$ such that $H^{\prime}$ contains a huge component and $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)$. By definition, $S$ is a maximal 2-node-connected subgraph of ^$\bm{\hat{}}{G}_{H}$ such that each node is part of a cycle of length at least 3.
If $S$ also contains a node $A$ corresponding to a $\mathcal{C}_{4}$ or a $\mathcal{C}_{5}$, then we apply Lemma 12 to obtain a cycle $K$ through $L$ and $A$ such that $K$ is incident to the vertices $u$ and $v$ of $C_{A}$. Now either Case (a) or Case (b) in the lemma statement holds.
If Case (a) holds, then there is a Hamiltonian Path $P$ from $u$ to $v$ in $G[V(C_{A})]$ and we set $H^{\prime}\coloneq(H\setminus E(C_{A}))\cup K\cup P$. Notice that $H^{\prime}$ is a canonical bridgeless 2-edge cover and has fewer components compared to $H$; in $H^{\prime}$ we have created a cycle using $K$ and $P$ that passes through some 2EC components (including $L$) and all the vertices of $C_{A}$ using the Hamiltonian path $P$. Now, observe that $|H^{\prime}|\leq|H|+|K|-1$, and as each component of $H$ has a credit of at least 1, the large component corresponding to $L$ has credit $2$, and the newly created huge component in $H^{\prime}$ receives a credit of 2, we have $\mathrm{cr}(H^{\prime})\leq\mathrm{cr}(H)-(|K|-1+2)+2=\mathrm{cr}(H)-|K|+1$. Therefore, $\mathrm{cost}(H^{\prime})=|H^{\prime}|+\mathrm{cr}(H^{\prime})\leq|H|+|K|-1+\mathrm{cr}(H)-|K|+1=|H|+\mathrm{cr}(H)=\mathrm{cost}(H)$.
Otherwise, Case (b) holds, and there exist a node $D$ which is not in the cycle $K$, and a set of edges $F$ disjoint to $K$, such that $|F|=|E(C_{A})|+|E(C_{D})|$ and $F\cup\{uv\}$ forms a 2-ECSS of $G[V(C_{A})\cup V(C_{D})])$. Here, we set $H^{\prime}\coloneq(H\setminus(E(C_{A})\cup E(C_{D})))\cup K\cup F$. Like before, $H^{\prime}$ is bridgeless and has fewer components compared to $H$ for the following reason. From the condition of Case (b), the vertices of $C_{A}$ and $C_{D}$ remain 2EC in $H^{\prime}$ with the edges of $F$ and an additional edge $uv$; in $H^{\prime}$ the work of the edge $uv$ is instead done by a $u-v$ path induced by the cycle $K$. Also, the cycle $K$ 2-edge connects the various 2EC components (including $L$) with the vertices of $C_{A}$ and $C_{D}$ with the help of $F$. Thus, no bridges are introduced. Furthermore, observe that $|H^{\prime}|\leq|H|+|K|$ as $|F|=|E(C_{A})|+|E(C_{D})|$, and since each component in $H$ has at least a credit of 1, the huge component corresponding to $L$ has credit $2$, $C_{D}$ is not a component of $H^{\prime}$ and was not on $K$, and the newly created huge component receives credit 2, we have $\mathrm{cr}(H^{\prime})\leq\mathrm{cr}(H)-(|K|-1+2+1)+2=\mathrm{cr}(H)-|K|$. Therefore, $\mathrm{cost}(H^{\prime})=|H^{\prime}|+\mathrm{cr}(H^{\prime})\leq|H|+|K|+\mathrm{cr}(H)-|K|=|H|+\mathrm{cr}(H)=\mathrm{cr}(H)$.
Hence, we can assume that $S$ does not contain a node corresponding to a $\mathcal{C}_{4}$ or a $\mathcal{C}_{5}$. Thus, all nodes of $S$ correspond to components that are either large, a $\mathcal{C}_{6}$ or a $\mathcal{C}_{7}$. Temporarily, we drop the parallel edges from the segment $S$ in the component graph ^$\bm{\hat{}}{G}_{H}$ keeping precisely one edge between adjacent nodes in $S$. Observe that $S$ continues to be 2-node connected even after dropping the parallel edges. Now, the smallest cycle through $L$ in $S$ in the component graph has length at least $3$ (as there are no parallel edges). Compute any cycle $K$ of length at least $3$ through $L$ in the segment $S$ and set $H^{\prime}\coloneq H\cup K$. $H^{\prime}$ is clearly a canonical bridgeless 2-edge cover of $G$ with fewer components. We have $|H^{\prime}|=|H|+|K|$ and since all components incident on the cycle $K$ have a credit of at least $\frac{3}{2}$ and $L$ has a credit of $2$, and they merge into a single huge component in $H^{\prime}$ which receives credit 2, we have $\mathrm{cr}(H^{\prime})\leq\mathrm{cr}(H)-(\frac{3}{2}(|K|-1)+2)+2=\mathrm{cr}(H)-\frac{3}{2}|K|+\frac{3}{2}\leq\mathrm{cr}(H)-|K|$, as $|K|\geq 3$ implies $\frac{|K|}{2}\geq\frac{3}{2}$. Combining the above facts, we have $\mathrm{cost}(H^{\prime})=|H^{\prime}|+\mathrm{cr}(H^{\prime})\leq|H|+|K|+\mathrm{cr}(H)-|K|\leq|H|+\mathrm{cr}(H)=\mathrm{cost}(H)$. ∎
6.3.1 Proof of Lemma 12
Only the proof of Lemma 12 is left. We will make use of the following results in our proof.
Proposition 13 (Theorem 1 in [25]).
Given a graph $G$ and an edge set $F\subseteq E(G)$ of constant size, we can find a cycle in $G$ that contains all edges of $F$ in polynomial time if one exists.
Proposition 14 (Lemma D.8 in [18]; reformulated).
Let $G$ be a graph and $H$ be a subgraph of $G$. Let $C$ and $C^{\prime}$ be two 2EC components of $H$ such that $C$ is a $\mathcal{C}_{5}$ and $C^{\prime}$ is a $\mathcal{C}_{4}$, and there is a $3$-matching between $C$ and $C^{\prime}$ in $G$. Then, for any two vertices $x,y\in V(C)$ there exists a set $F$ of $9$ edges in $G$ such that $F\cup\{xy\}$ induces a 2EC spanning subgraph of $V(C)\cup V(C^{\prime})$ in $G$.
Furthermore, the following simple lemma showing the existence of certain $3$-matchings within a non-trivial segment will also be used in the proof.
Lemma 15.
Let $S$ be a non-trivial segment of ^$\bm{\hat{}}{G}_{H}$ and $A$ be a node in $S$. Then, there exists a matching of size 3 in $G$ between $V(C_{A})$ and $V(S)\setminus V(C_{A})$.
Proof.
If $A$ is not a cut node in ^$\bm{\hat{}}{G}_{H}$, then, by the definition of a segment, there are no edges between $V(C_{A})$ and the vertices of components which are not contained in $S$. Thus, in this case the claim follows from Lemma 2 using $(V(C_{A}),V(G)\setminus V(C_{A}))$ as the partition of $V(G)$. Otherwise, $A$ is a cut node in ^$\bm{\hat{}}{G}_{H}$, which implies $C_{A}$ is a separator of $G$ with the vertices of $V(S)\setminus V(C_{A})$ on one side of the separator (possibly along with some other vertices of $V(G)$). Consequently, in this case, our claim follows by applying Lemma 2 on a partition of $V(G)$ into two parts, where one part contains the vertices of $V(S)\setminus V(C_{A})$ (one side of the separator mentioned above) and the other part contains the vertices from $V(C_{A})$ such that the only edges crossing the two parts are between $V(S)\setminus V(C_{A})$ and $V(C_{A})$. ∎
We are now ready to prove Lemma 12.
Proof of Lemma 12.
We are given a non-trivial segment $S$ of $G_{H}$ that contains a huge node $L$ and a node $A$ such that $C_{A}$ is either a $\mathcal{C}_{4}$ or a $\mathcal{C}_{5}$. We need to construct in polynomial time a cycle $K$ in $G_{H}$ through $L$ and $A$, such that the two cycle edges incident on $C_{A}$ are incident on two vertices $u$ and $v$ in $C_{A}$ such that either (a) we can find a $u-v$ Hamiltonian path in $G[V(C_{A})]$ or (b) we can find a node $D$ outside the cycle $K$ and a set of edges $F$ disjoint to $E(K)$ such that $|F|=|E(C_{A})|+|E(C_{D})|$ and $F\cup\{uv\}$ forms a 2-ECSS of $G[V(C_{A})\cup V(C_{D})]$.
In the proof we will show the existence of the cycle $K$ and a desired Hamiltonian path or a desired $D$ and $F$ with $|F|=9$. The polynomial-time construction for these objects then easily follow; $K$ from the algorithm in Proposition 13 by feeding it all possible $4$ edges, $2$ incident on $L$ and $2$ incident on $A$, and then the Hamiltonian path or $F$ can be found by brute-force enumeration over a constant domain.
Let the cycle $C_{A}$ be labelled as $v_{1}-v_{2}-v_{3}-v_{4}-v_{1}$ if it is a $\mathcal{C}_{4}$ or $v_{1}-v_{2}-v_{3}-v_{4}-v_{5}-v_{1}$ if is is a $\mathcal{C}_{5}$. Since $S$ is a segment, it is 2-node-connected and, thus, there exist two internally node-disjoint paths $P_{1}$ and $P_{2}$ in $S$ between $L$ and $A$. Moreover, Lemma 15 implies that at least three vertices of $V(C_{A})$ have incident edges in $G$ from the vertices in $S$. We consider two exhaustive cases based on which set of vertices of $C_{A}$ has such incident edges: Either (Case 1) each of $v_{1},v_{2},v_{3}$ or (Case 2) each of $v_{1},v_{2},v_{4}$ has an incident edge from the vertices of $S$ that are not in $C_{A}$. Observe up to symmetry these are the only cases, and for $C_{A}$ being a $\mathcal{C}_{4}$ considering only Case 1 is sufficient. Thus, in Case 2, we will assume $C_{A}$ is a $\mathcal{C}_{5}$.
Case 1: We have two subcases.
Case 1.1: There exists paths $P_{1}$ and $P_{2}$ that are incident to vertices $u$ and $v$ of $C_{A}$ such that there exists a Hamiltonian Path between $u$ and $v$ in $G[V(C_{A})]$. Then, $K=P_{1}\cup P_{2}$ forms a cycle in ^$\bm{\hat{}}{G}_{H}$ through $A$ and $L$ as $P_{1}$ and $P_{2}$ are internally node-disjoint, which satisfies the requirements of (a).
Case 1.2: Otherwise, if we are not in Case 1.1, there must exist paths $P_{1}$ and $P_{2}$ such that $P_{1},P_{2}$ are incident to $V(C_{A})$ at vertices $u_{1},u_{2}\in V(C_{A})$, respectively ($u_{1}=u_{2}$ is possible), and there exist $w_{i}\in\{v_{1},v_{2},v_{3}\}$ adjacent to $u_{i}$ in $C_{A}$ for $i\in\{1,2\}$. This observation uses that $C_{A}$ is a cycle of at most 5 vertices. Consider two subcases:
[…]
Case 1.2.2: Since we are in Case 1.2 and not in Case 1.2.1, there only exist paths $P_{1},P_{2}$ and $w_{1},w_{2}\in\{v_{1},v_{2},v_{3}\}$ such that $u_{1}\neq u_{2}$ and $w_{1}\neq w_{2}$. Since there is no Hamiltonian path between $u_{1}$ and $u_{2}$ (Case 1.1), it must be that $u_{1}$ and $u_{2}$ are not adjacent in the cycle $C_{A}$. Additionally, $C_{A}$ must be a $\mathcal{C}_{5}$ ($C_{A}=\mathcal{C}_{4}$, $u_{1}\neq u_{2}$, and $u_{1}$ not adjacent to $u_{2}$ implies they are diagonally opposite, and we can set $w_{1}=w_{2}\in\{v_{1},v_{2},v_{3}\}$). Furthermore, we can set $w_{1}$ and $w_{2}$ such that they are adjacent and in $\{v_{1},v_{2},v_{3}\}$ (this can be seen from the case we are in: $u_{1}\neq u_{2}$, $u_{1}$ and $u_{2}$ are not adjacent, $C_{A}=\mathcal{C}_{5}$ and it is not possible to set $w_{1}=w_{2}\in\{v_{1},v_{2},v_{3}\}$). Let $A_{i}\neq A$ be the node in $S$ that is adjacent to $w_{i}$, for $i\in\{1,2\}$. Since $S$ is 2-node connected, there must be a path $P^{\prime}_{i}$ in $S$ from $A_{i}$ to $L$ which does not contain $A$, for $i\in\{1,2\}$.
[…]
Case 2: The vertices $v_{1}$, $v_{3}$, and $v_{4}$ have incident edges to nodes in $S$ other than $A$. As discussed above, this case can only apply if $A$ corresponds to a $\mathcal{C}_{5}$. Further, note that neither $v_{2}$ nor $v_{5}$ can have an incident edge to a node in $S$ other than $A$, as otherwise we are also in Case 1 by simply relabelling the $v_{i}$’s cyclically so that $v_{1},v_{2},v_{3}$ have such incident edges. We have two subcases based on whether or not $v_{2}$ or $v_{5}$ has an incident edge to a component outside of segment $S$.
[…]
Case 2.2: There is an edge between $v_{2}$ (or $v_{5}$ by symmetry) and a node $D$ of ^$\bm{\hat{}}{G}_{H}$ which is a trivial segment $S^{\prime}$ or lies in a non-trivial segment $S^{\prime}$, and $S^{\prime}\neq S$. Note that $A$ must be a cut node in ^$\bm{\hat{}}{G}_{H}$ in this case. Observe that $|V(S)|-|V(C_{A})|\geq 10$, because $S$ contains the node $L$ ($C_{L}$ is huge with 10 vertices) other than $A$. If $|V(S^{\prime})\cup V(C_{A})|\geq 10$, Lemma 3 guarantees the existence of a $4$-matching $M_{4}$ between $V(S)\setminus V(C_{A})$ and $V(C_{A})$ (as $C_{A}$ is a separator, $V(G)$ can be partitioned in 2 parts with $V(C_{A})\cup V(S^{\prime})$ on one side and $V(S)\setminus V(C_{A})$ on the other, so that the only edges crossing are between $V(C_{A})$ and $V(S)\setminus V(C_{A})$). But this means that we are actually in Case 1 as some three consecutive vertices of $C_{A}$ must be incident to the other nodes of $S$.
[…]
Appendix A Reduction to Structured Graphs
This section is dedicated to the proof of Definition 5, which we first restate.
[…]
We assume throughout this section that $\alpha\geq\frac{5}{4}$ and $\varepsilon\in(0,\frac{1}{24}]$. Let $\mathsf{ALG}$ be a polynomial-time algorithm that, given an $(\alpha,\varepsilon)$-structured graph $G^{\prime}=(V^{\prime},E^{\prime})$, returns an $\alpha$-approximate 2EC spanning subgraph $\mathsf{ALG}(G^{\prime})\subseteq E^{\prime}$ of $G^{\prime}$. In the following, we define a reduction algorithm which recursively modifies and partitions the input graph $G$ until we obtain $(\alpha,\varepsilon)$-structured graphs, then solves these with $\mathsf{ALG}$, and finally puts the resulting solutions together to an overall solution for $G$. In particular, our reduction checks for structures that are forbidden in $(\alpha,\varepsilon)$-structured graphs. These are graphs of constant size (w.r.t. $\varepsilon$), parallel edges and self loops, 1-vertex cuts, small $\alpha$-contractible subgraphs (w.r.t. $\varepsilon$), irrelevant edges, non-isolating 2-vertex cuts and large 3-vertex cuts (cf. Definition 5). For all structures except large 3-vertex cuts we can use the reduction presented in [18]. The main contribution of this section is to show how to reduce to graphs without large 3-vertex cuts, given that all other substructures have already been removed.
A.1 Definitions and the Algorithm
We will heavily use the following two facts on 2EC graphs with contracted components. We will not refer to them explicitly.
Fact 16.
Let $G$ be a 2EC graph and $W\subseteq V(G)$. Then $G|W$ is 2EC.
Fact 17.
Let $H$ be a 2EC subgraph of a 2EC graph $G$, and $S$ and $S^{\prime}$ be 2EC spanning subgraphs of $H$ and $G|H$, respectively. Then $S\cup S^{\prime}$ is a 2EC spanning subgraph of $G$, where we interpret $S^{\prime}$ as edges of $G$ after decontracting $V(H)$.
Next, we classify an optimal solution $\mathsf{OPT}(G)$ with respect to a large 3-vertex cut.
Definition 13 (Solution types).
Let $H$ be a graph and $u,v,w\in V(H)$. Let $H^{\prime}$ be obtained by contracting each 2EC component $C$ of $H$ into a single super-node $C$, and let $C(x)$ be the corresponding super-node of the 2EC component of $H$ that contains $x\in V(H)$. We define the following types of $H^{\prime}$ with respect to $\{u,v,w\}$:
Type $\mathsf{A}$:
$H^{\prime}$ is composed of a single super-node $H^{\prime}=C(u)=C(v)=C(w)$.
Type $\mathsf{B1}$:
$H^{\prime}$ is a $C_{1}-C_{k}$ path of length at least $1$ such that $u,v,w\in C_{1}\cup C_{k}$ and |$\lvert\{u,v,w\}\cap C_{i}\rvert\leq 2$ for $i\in\{1,k\}$. We assume w.l.o.g. that $C_{1}=C(u)=C(v)$.
Type $\mathsf{B2}$:
$H^{\prime}$ is composed of two isolated super-nodes $C_{1}$ and $C_{2}$, where |$\lvert\{u,v,w\}\cap C_{i}\rvert\leq 2$ for $i\in\{1,2\}$. We assume w.l.o.g. that $C_{1}=C(u)=C(v)$.
Type $\mathsf{C1}$:
$H^{\prime}$ is a tree and $u$, $v$, and $w$ are in distinct super-nodes.
Type $\mathsf{C2}$:
$H^{\prime}$ is composed of a $C_{1}-C_{k}$ path of length at least $1$ and an isolated super-node $C_{\ell}$ such that $\lvert C_{i}\cap\{u,v,w\}\rvert=1$ for $i\in\{1,k,\ell\}$.
Type $\mathsf{C3}$:
$H^{\prime}$ is composed of the three isolated super-nodes $C(u),C(v)$ and $C(w)$.
Let $\mathcal{T}\coloneq\{\mathsf{A},\mathsf{B1},\mathsf{B2},\mathsf{C1},\mathsf{C2},\mathsf{C3}\}$ be the set of 3-vertex cut optimal solution types.
Definition 14 (Order on types).
$\mathsf{A}\succ\mathsf{B1}\succ\mathsf{B2}\succ\mathsf{C1}\succ\mathsf{C2}\succ\mathsf{C3}$.
The following proposition lists compatible solution types of both sides of a 3-vertex cut. The proof is straightforward and, thus, omitted.
Proposition 18 (feasible type combinations).
Let $\{u,v,w\}$ be a large 3-vertex cut in a graph $G$, and $(V_{1},V_{2})$ be a partition of $V\setminus\{u,v,w\}$ such that $7\leq\lvert V_{1}\rvert\leq\lvert V_{2}\rvert$. Let $G_{1}=G[V_{1}\cup\{u,v,w\}]$, $G_{2}=G[V_{1}\cup\{u,v,w\}]\setminus\{uv,vw,uw\}$ and, for a fixed optimal solution $\mathsf{OPT}(G)$, let $\mathsf{OPT}_{i}=\mathsf{OPT}(G)\cap E(G_{i})$ for $i\in\{1,2\}$. Then, with respect to $\{u,v,w\}$, if
  1. (a)
    $\mathsf{OPT}_{1}$ is of type $\mathsf{A}$, $\mathsf{OPT}_{2}$ must be of type $t\in\{\mathsf{A},\mathsf{B1},\mathsf{B2},\mathsf{C1},\mathsf{C2},\mathsf{C3}\}=\mathcal{T}$.
  2. (b)
    $\mathsf{OPT}_{1}$ is of type $\mathsf{B1}$, $\mathsf{OPT}_{2}$ must be of type $t\in\{\mathsf{A},\mathsf{B1},\mathsf{B2},\mathsf{C1},\mathsf{C2}\}$.
  3. (c)
    $\mathsf{OPT}_{1}$ is of type $\mathsf{B2}$, $\mathsf{OPT}_{2}$ must be of type $t\in\{\mathsf{A},\mathsf{B1},\mathsf{B2}\}$.
  4. (d)
    $\mathsf{OPT}_{1}$ is of type $\mathsf{C1}$, $\mathsf{OPT}_{2}$ must be of type $t\in\{\mathsf{A},\mathsf{B1},\mathsf{C1}\}$.
  5. (e)
    $\mathsf{OPT}_{1}$ is of type $\mathsf{C2}$, $\mathsf{OPT}_{2}$ must be of type $t\in\{\mathsf{A},\mathsf{B1}\}$.
  6. (f)
    $\mathsf{OPT}_{1}$ is of type $\mathsf{C3}$, $\mathsf{OPT}_{2}$ must be of type $\mathsf{A}$.
Observe that according to the above definition we have $\lvert\mathsf{OPT}(G)\rvert=\lvert\mathsf{OPT}_{1}\rvert+\lvert\mathsf{OPT}_{2}\rvert$.
We are now ready to state our main algorithm $\mathsf{RED}$ (Algorithm 1). The reductions applied in Algorithms 1, 1, 1, 1 and 1 are identical to the reductions given in [18]. When Algorithm 1 reaches Algorithm 1 it calls Algorithm 2 to handle a large 3-vertex cut $\{u,v,w\}$. Whenever the reduction returns edges $K$ of contracted graphs $G|H$, i.e., $K=\mathsf{RED}(G|H)$, we interpret these as the corresponding edges after decontracting $H$, e.g., in Algorithm 1 in Algorithm 1 or in Algorithm 2 in Algorithm 2.
Hence, from now on let us focus on Algorithm 2. In Section A.2 we prove that the claimed sets of edges $F$ in Algorithm 2 exist and are of constant size. The computations done in Algorithm 2 can be done as explained in [18]. That is, we first compute via enumeration minimum-size subgraphs $\overline{\mathsf{OPT}}_{1}^{t}$ of $G_{1}$ of every type $t\in\mathcal{T}$ if such a solution exists, because $G_{1}$ is of constant size in this case. If $\overline{\mathsf{OPT}}_{1}^{t}$ does not exist, $\mathsf{OPT}_{1}^{t}$ also cannot exist. Otherwise, that is, $\overline{\mathsf{OPT}}_{1}^{t}$ exists, we check whether $G_{2}$ is a graph that admits a solution that is compatible with $\overline{\mathsf{OPT}}_{1}^{t}$ according to Proposition 18, and, if that is the case, we set $\mathsf{OPT}_{1}^{t}\coloneq\overline{\mathsf{OPT}}_{1}^{t}$. Finally note that if the algorithm reaches Algorithm 1 in Algorithm 1, then $G$ must be $(\alpha,\varepsilon)$-structured.
Input: 2EC graph $G=(V,E)$.
1 if $V(G)\leq\frac{4}{\varepsilon}$ then
2       compute $\mathsf{OPT}(G)$ via enumeration and return $\mathsf{OPT}(G)$
3if $G$ has a 1-vertex cut $\{v\}$  then
4       let $(V_{1},V_{2})$, be a partition of $V\setminus\{v\}$ such that $V_{1}\neq\emptyset\neq V_{2}$ and there are no edges between $V_{1}$ and $V_{2}$.
5       return $\mathsf{RED}(G[V_{1}\cup\{v\}])\cup\mathsf{RED}(G[V_{2}\cup\{v\}])$.
6
7if $G$ contains a self loop or a parallel edge $e$  then
8       return $\mathsf{RED}(G\setminus\{e\})$.
9
10if $G$ contains an $\alpha$-contractible subgraph $H$ with $V(H)\leq\frac{4}{\varepsilon}$  then
11       return $H\cup\mathsf{RED}(G\mid H)$.
12if $G$ contains an irrelevant edge $e$  then
13       return $\mathsf{RED}(G\setminus\{e\})$.
14
15if $G$ contains a non-isolating 2-vertex cut $\{u,v\}$  then
16       Execute Lines 12 to 25 of Algorithm 1 in [18] for $\{u,v\}$.
17
18if $G$ contains a large 3-vertex cut $\{u,v,w\}$  then
19       Execute Algorithm 2 for $\{u,v,w\}$.
return $\mathsf{ALG}(G)$.
Algorithm 1 $\mathsf{RED}(G)$: reduction to $(\alpha,\varepsilon)$-structured graphs
Input: A 2EC graph $G=(V,E)$ that does not contain cut nodes, non-isolating 2-vertex cuts, or $\alpha$-contractible subgraphs with at most $\frac{4}{\varepsilon}$ vertices. A large 3-vertex cut $\{u,v,w\}$ in $G$.
1 Let $(V_{1},V_{2})$ be a partition of $V\setminus\{u,v,w\}$ such that $7\leq\lvert V_{1}\rvert\leq\lvert V_{2}\rvert$ and there are no edges between $V_{1}$ and $V_{2}$ in $G$.
2 Let $G_{1}\coloneq G[V_{1}\cup\{u,v,w\}]$ and $G_{2}\coloneq G[V_{2}\cup\{u,v,w\}]\setminus\{uv,vw,uw\}$.
3 Let $G^{\prime}_{i}\coloneq G_{i}|\{u,v,w\}$ for $i\in\{1,2\}$.
4 if $\lvert V_{1}\rvert>\frac{2}{\varepsilon}-4$ then
5       Let $H^{\prime}_{i}\coloneq\mathsf{RED}(G^{\prime}_{i})$ for $i\in\{1,2\}$.
6       Let $F^{\prime}\subseteq E$ be a minimum-size edge set such that $H^{\prime}\coloneq H^{\prime}_{1}\cup H^{\prime}_{2}\cup F^{\prime}$ is 2EC. return $H^{\prime}$
7
8else
9       Let $\mathsf{OPT}_{1}^{t}$ be the minimum-size subgraphs of $G_{1}$ for every type $t\in\mathcal{T}$ (w.r.t. $\{u,v,w\}$ and $G_{1}$) that belong to some 2EC spanning subgraph of $G$, if exists. Let $\mathsf{OPT}_{1}^{\min}$ be the existing subgraph of minimum size among all types in $\mathcal{T}$, where ties are broken according to Definition 14 (that is, $t^{\prime}\in\mathcal{T}$ is preferred over $t\in\mathcal{T}$ if $t^{\prime}\succ t$), and $t^{\min}\in\mathcal{T}$ the corresponding type. Let $\mathsf{opt}_{1}^{t}\coloneq|\mathsf{OPT}_{1}^{t}|$ for every $t\in\mathcal{T}$ and $\mathsf{opt}_{1}^{\min}\coloneq|\mathsf{OPT}_{1}^{\min}|$.
10       if $\mathsf{OPT}_{1}^{\mathsf{B1}}$ exists and $\mathsf{opt}_{1}^{\mathsf{B1}}\leq\mathsf{opt}_{1}^{\min}+1$ then // also if $t^{\min}=\mathsf{B1}$
11             Let $G_{2}^{\mathsf{B1}}\coloneq G_{2}^{\prime}$ and $H_{2}^{\mathsf{B1}}\coloneq\mathsf{RED}(G_{2}^{\mathsf{B1}})$.
12             Let $F^{\mathsf{B1}}\subseteq E$ be a min-size edge set s.t. $H^{\mathsf{B1}}\coloneq\mathsf{OPT}_{1}^{\mathsf{B1}}\cup H_{2}^{\mathsf{B1}}\cup F^{\mathsf{B1}}$ is 2EC. return $H^{\mathsf{B1}}$.
13
14      else if $t^{\min}=\mathsf{B2}$ then
15             Let $G_{2}^{\mathsf{B2}}\coloneq G_{2}^{\prime}$ and $H_{2}^{\mathsf{B2}}\coloneq\mathsf{RED}(G_{2}^{\mathsf{B2}})$.
16             Let $F^{\mathsf{B2}}\subseteq E$ be a min-size edge set s.t. $H^{\mathsf{B2}}\coloneq\mathsf{OPT}_{1}^{\mathsf{B2}}\cup H_{2}^{\mathsf{B2}}\cup F^{\mathsf{B2}}$ is 2EC. return $H^{\mathsf{B2}}$.
17
18      else if $t^{\min}=\mathsf{C1}$ then
19             Let $G_{2}^{\mathsf{C1}}\coloneq G_{2}^{\prime}$ and $H_{2}^{\mathsf{C1}}\coloneq\mathsf{RED}(G_{2}^{\mathsf{C1}})$.
20             Let $F^{\mathsf{C1}}\subseteq E$ be a min-size edge set s.t. $H^{\mathsf{C1}}\coloneq\mathsf{OPT}_{1}^{\mathsf{C1}}\cup H_{2}^{\mathsf{C1}}\cup F^{\mathsf{C1}}$ is 2EC. return $H^{\mathsf{C1}}$.
21
22      else if $t^{\min}=\mathsf{C2}$ then
23             if Every $\mathsf{OPT}_{1}^{\mathsf{C2}}$ solution contains a $C(u)-C(v)$ path. then
24                   Let $G_{2}^{\mathsf{C2}}\coloneq(V(G_{2})\cup\{y\},E(G_{2})\cup\{uy,vy,vw\})$.
25                   Let $H_{2}^{\mathsf{C2}}\coloneq\mathsf{RED}(G_{2}^{\mathsf{C2}})\setminus\{uy,vy,vw\}$.
26                   Let $F^{\mathsf{C2}}\subseteq E$ be a min-size edge set s.t. $H^{\mathsf{C2}}\coloneq\mathsf{OPT}_{1}^{\mathsf{C2}}\cup H_{2}^{\mathsf{C2}}\cup F^{\mathsf{C2}}$ is 2EC. return $H^{\mathsf{C2}}$.
27
28            else if Every $\mathsf{OPT}_{1}^{\mathsf{C2}}$ solution contains either a $C(u)-C(v)$ or a $C(v)-C(w)$ path. then
29                   Let $G_{2}^{\mathsf{C2}}\coloneq(V(G_{2})\cup\{y,z\},E(G_{2})\cup\{uy,vz,zy,wy\})$.
30                   Let $H_{2}^{\mathsf{C2}}\coloneq\mathsf{RED}(G_{2}^{\mathsf{C2}})\setminus\{uy,vz,zy,wy\}$.
31                   Let $F^{\mathsf{C2}}\subseteq E$ be a min-size edge set and $\mathsf{OPT}_{1}^{\mathsf{C2}}$ be s.t. $H^{\mathsf{C2}}\coloneq\mathsf{OPT}_{1}^{\mathsf{C2}}\cup H_{2}^{\mathsf{C2}}\cup F^{\mathsf{C2}}$ is 2EC. return $H^{\mathsf{C2}}$.
32
33            else // all paths are possible between $C(u)$, $C(v)$, and $C(w)$
34                   Let $G_{2}^{\mathsf{C2}}\coloneq(V(G_{2})\cup\{y\},E(G_{2})\cup\{uy,vy,wy\})$.
35                   Let $H_{2}^{\mathsf{C2}}\coloneq\mathsf{RED}(G_{2}^{\mathsf{C2}})\setminus\{uy,vy,wy\}$.
36                   Let $F^{\mathsf{C2}}\subseteq E$ be a min-size edge set and $\mathsf{OPT}_{1}^{\mathsf{C2}}$ be s.t. $H^{\mathsf{C2}}\coloneq\mathsf{OPT}_{1}^{\mathsf{C2}}\cup H_{2}^{\mathsf{C2}}\cup F^{\mathsf{C2}}$ is 2EC. return $H^{\mathsf{C2}}$.
37
38
39      else if $t^{\min}=\mathsf{C3}$ then
40             Let $\mathsf{OPT}_{1}^{\mathsf{C3}}$ be a $\mathsf{C3}$ solution such that there is an edge $e^{\mathsf{C3}}_{uv}$ in $G_{1}$ between $C(u)$ and $C(v)$ and there is an edge $e^{\mathsf{C3}}_{vw}$ in $G_{1}$ between $C(v)$ and $C(w)$.
41             Let $G_{2}^{\mathsf{C3}}\coloneq(V(G_{2}),E(G_{2})\cup\{uv,uv,vw,vw\})$ and $H_{2}^{\mathsf{C3}}\coloneq\mathsf{RED}(G_{2}^{\mathsf{C3}})\setminus\{uv,uv,vw,vw\}$.
42             Let $F^{\mathsf{C3}}\subseteq E$ be a min-size edge set s.t. $H^{\mathsf{C3}}\coloneq\mathsf{OPT}_{1}^{\mathsf{C3}}\cup H_{2}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ is 2EC. return $H^{\mathsf{C3}}$.
43
44
Algorithm 2 Remove a large 3-vertex cut
A.2 Helper Lemmas
Proposition 19.
If Algorithm 1 executes Algorithm 2, then $G$ contains no non-isolating 2-vertex cuts, self loops, parallel edges, irrelevant edges, or $\alpha$-contractable subgraphs with at most $\frac{4}{\varepsilon}$ vertices.
Lemma 20.
Let $G$ be the graph when Algorithm 1 executes Algorithm 2, $\{u,v,w\}$ be a large 3-vertex cut of $G$, $(V_{1},V_{2})$ be partition of $V\setminus\{u,v\}$ such that $V_{1}\neq\emptyset\neq V_{2}$ and there are no edges between $V_{1}$ and $V_{2}$, and $H$ be a 2EC spanning subgraph of $G$. Let $G_{i}=G[V_{i}\cup\{u,v,w\}]$ and $H_{i}=E(G_{i})\cap H$ for $i\in\{1,2\}$. Then, $H_{1}$ and $H_{2}$ are of a type in $\mathcal{T}$ with respect to  $\{u,v,w\}$ and $G_{1}$ and $G_{2}$, respectively.
Proof.
Let $i\in\{1,2\}$. First, observe that $H_{i}$ consist of either one, two or three connected components, and each component contains at least one of $\{u,v,w\}$, as otherwise $H$ would be disconnected. Let $H_{i}^{\prime}$ be obtained by contracting each 2EC component $C$ of $H_{i}$ into a single super-node $C$, and let $C_{i}(x)$ be the corresponding super-node of the 2EC component of $H$ that contains $x\in V(H_{i})$.
First, assume that $H_{i}$ is a single connected component. In this case $H_{i}^{\prime}$ is a tree. If $H_{i}^{\prime}$ is composed of a single node, $H_{i}$ is of type $\mathsf{A}$. Otherwise, let $F$ be the set of edges of $H_{i}^{\prime}$ that are on any simple path between $C_{i}(u)$, $C_{i}(v)$, and $C_{i}(w)$. If $H_{i}^{\prime}$ contains an edge $e\in E(H_{i}^{\prime})\setminus F$, then $e$ cannot belong to any cycle of $H$, because it does not belong to a cycle in $H_{i}$ and is not on a simple path between any two vertices of the cut $\{u,v,w\}$, contradicting that $H$ is 2EC. Thus, $H_{i}$ is of type $\mathsf{C1}$ if $C_{i}(u)$, $C_{i}(v)$, and $C_{i}(w)$ are distinct, and of type $\mathsf{B1}$ otherwise.
Second, assume that $H_{i}$ is composed of two connected components. Let $H_{i}^{\prime}(v)$ and $H_{i}^{\prime}(w)$ denote the two trees in $H_{i}^{\prime}$ such that w.l.o.g. $u$ and $v$ are in $H_{i}^{\prime}(v)$ and $w$ is in $H_{i}^{\prime}(w)$. First note that $H_{i}^{\prime}(w)$ must be a single node. Otherwise, there is an edge $e$ in $H_{i}^{\prime}(w)$ that cannot be part of any cycle in $H$, contradicting that $H$ is 2EC. If $H_{i}^{\prime}(v)$ is also a single node, then $H_{i}$ is of type $\mathsf{B2}$. Otherwise, let $F$ be the set of edges of $H_{i}^{\prime}(v)$ on the simple path between $C_{i}(u)$ and $C_{i}(v)$. If $H_{i}^{\prime}(v)$ contains an edge $e\in E(H_{i}^{\prime}(v))\setminus F$, then $e$ cannot belong to any cycle of $H$, contradicting that $H$ is 2EC. Thus, $H_{i}$ is of type $\mathsf{C2}$.
Finally, assume that $H_{i}$ is composed of three connected components. Let $H_{i}^{\prime}(u)$, $H_{i}^{\prime}(v)$, and $H_{i}^{\prime}(w)$ denote the trees of $H_{i}^{\prime}$ that contain $u$, $v$, and $w$, respectively. Using the same argument as before, we can show that each tree is composed of a single node, and, thus, $H_{i}$ is of type $\mathsf{C3}$. ∎
Lemma 21.
In Algorithm 2 it holds that $\mathsf{opt}_{1}^{\min}\geq 8$.
Proof.
Consider any feasible 2-ECSS solution $H_{1}$ for $G_{1}$, and let $H^{\prime}_{1}=H_{1}\setminus\{u,v,w\}$ be the corresponding solution for $G_{1}|\{u,v,w,\}$. Since $G_{1}|\{u,v,w,\}$ contains at least $8$ vertices and $H^{\prime}_{1}$ is a feasible 2-ECSS solution, $H^{\prime}_{1}$ must contain at least $8$ edges. Thus $|H_{1}|\geq|H^{\prime}_{1}|\geq 8$. ∎
Lemma 22.
If Algorithm 2 reaches Algorithm 2 and $\mathsf{OPT}_{1}^{\mathsf{A}}$ exists, then $\mathsf{opt}_{1}^{\mathsf{A}}\geq\alpha\cdot\mathsf{opt}_{1}^{\min}+1\geq\mathsf{opt}_{1}^{\min}+3$.
Proof.
If $\mathsf{OPT}_{1}^{\mathsf{A}}$ exists and $\mathsf{opt}_{1}^{\mathsf{A}}\leq\alpha\cdot\mathsf{opt}_{1}^{\min}$ then $\mathsf{OPT}_{1}^{\mathsf{A}}$ is 2EC and $\alpha$-contractible. However, since we assume that we reached Algorithm 2, we have that $|V_{1}|\leq\frac{2}{\varepsilon}-4\leq\frac{4}{\varepsilon}$. Thus, this is a contradiction to Proposition 19, as we would have contracted $G_{1}$ in Algorithm 1 earlier. Thus, it must be that $\mathsf{opt}_{1}^{\mathsf{A}}\geq\alpha\cdot\mathsf{opt}_{1}^{\min}+1$. Then, $\alpha\cdot\mathsf{opt}_{1}^{\min}\geq\mathsf{opt}_{1}^{\min}+2$, since $\alpha\geq\frac{5}{4}$ and $\mathsf{opt}_{1}^{\min}\geq 8$ by Lemma 21. ∎
We have an immediate corollary from this lemma, which justifies that we do not need to consider the case $t^{\min}=\mathsf{A}$ in Algorithm 2.
Corollary 23.
If Algorithm 2 reaches Algorithm 2, then $t^{\min}\neq\mathsf{A}$.
Lemma 24.
Let $G=(V,E)$ be a (multi-)graph that does not contain cut nodes or non-isolating 2-vertex cuts, and let $\{u,v,w\}$ be a large 3-vertex cut of $G$. Let $(V_{1},V_{2})$ be a partition of $V\setminus\{u,v,w\}$ such that $7\leq\lvert V_{1}\rvert\leq\lvert V_{2}\rvert$ and there are no edges between $V_{1}$ and $V_{2}$ in $G$, and let $G_{1}\coloneq G[V_{i}\cup\{u,v,w\}]$ and $G_{2}\coloneq G[V_{i}\cup\{u,v,w\}]\setminus\{uv,vw,uw\}$. For $i\in\{1,2\}$, let $H_{i}\subseteq E(G_{i})$, let $H_{i}(z)$ be the connected component of $H_{i}$ that contains $z\in V(H_{i})$ and let $H_{i}$ be such that $V(H_{i})=V(H_{i}(u))\cup V(H_{i}(v))\cup V(H_{i}(w))$, i.e., each vertex $y\in V(H_{i})$ is connected to $u$, $v$, or $w$ in $H_{i}$. Then, the following holds.
  1. 1)
    For $x\in\{u,v,w\}$ and $i\in\{1,2\}$, if $H_{i}(x)\neq H_{i}$, then there exists an edge $e\in E(G_{i})\setminus H_{i}$ between the vertices of $H_{i}(x)$ and the vertices of $H_{i}\setminus H_{i}(x)$.
  2. 2)
    For $i\in\{1,2\}$, if $H_{i}(u)$, $H_{i}(v)$ and $H_{i}(w)$ are all pairwise vertex-disjoint, then there exist edges $F\subseteq E(G_{i})\setminus H_{i}$ such that $|F|\geq 2$ and $F\subseteq\{e_{uv},e_{vw},e_{uw}\}$ where $e_{xy}$ is between the vertices of $H_{i}(x)$ and the vertices of $H_{i}(y)$.
Proof.
We prove the first claim, and note that the second claim follows from applying the first claim twice. Let $i\in\{1,2\}$, $x\in\{u,v,w\}$ and $H_{i}(x)\neq H_{i}$. If $H_{i}(x)$ is composed of the single vertex $x$ such that there is no edge between $x$ and $H_{i}\setminus H_{i}(x)$, then $\{u,v,w\}\setminus\{x\}$ must be a non-isolating two-vertex cut in $G$, a contradiction. Otherwise, that is, $H_{i}(x)$ consists of at least two vertices, and there is no edge between $H_{i}(x)$ and $H_{i}\setminus H_{i}(x)$ then $x$ is a cut node in $G$ (separating $V(H_{i}(x))\setminus\{x\}$ from $V\setminus V(H_{i}(x))$), a contradiction. ∎
Lemma 25.
Let $H\subseteq E$ be such that $H$ is connected and contains a spanning subgraph of $G$, and let $H_{i}\coloneq H\cap E(G_{i})$ for $i\in\{1,2\}$. If for all $i\in\{1,2\}$ every edge $e\in H_{i}$ is part of a 2EC component of $H_{i}$ or lies on an $x-y$ path in $H_{i}$ where $x,y\in\{u,v,w\}$ such that $x\neq y$ and there exists an $x-y$ path in $H_{j}$ where $j=1$ if $i=2$ and $j=2$ if $i=1$, then $H$ is a 2EC spanning subgraph of $G$.
Proof.
We show that every edge $e\in H$ is part of a 2EC component of $H$, which proves the lemma because $H$ is connected and spanning. If for all $i\in\{1,2\}$ every edge $e\in H_{i}$ is part of a 2EC component of $H_{i}$, it is also part of a 2EC component of $H$. Otherwise, for all $i\in\{1,2\}$, if $e$ lies on an $x-y$ path in $H_{i}$ where $x,y\in\{u,v,w\}$ such that $x\neq y$ and there exists an $x-y$ path in $H_{j}$ where $j=1$ if $i=2$ and $j=2$ if $i=1$, then $e$ is on a cycle of $H$ as both paths are edge-disjoint. ∎
Lemma 26.
If the condition in Algorithm 2 in Algorithm 2 applies and both, $H^{\prime}_{1}$ and $H^{\prime}_{2}$ are 2EC spanning subgraphs of $G^{\prime}_{1}$ and $G^{\prime}_{2}$, respectively, then there exist edges $F^{\prime}$ with $|F^{\prime}|\leq 4$ such that $H^{\prime}$ is a 2EC spanning subgraph of $G$.
Proof.
Let $i\in\{1,2\}$, and let $H^{\prime}_{i}(u)$, $H^{\prime}_{i}(v)$, and $H^{\prime}_{i}(w)$ be the connected components of $H^{\prime}_{i}$ after decontracting $\{u,v,w\}$ that contain $u$, $v$ and $w$, respectively. By Lemma 24, there exist edges $F_{i}\subseteq E(G_{i})$ such that $|F_{i}|\leq 2$ and $H^{\prime}_{i}\cup F_{i}$ is connected in $G_{i}$. Since $H^{\prime}_{i}$ is spanning in $G_{i}$, setting $F\coloneq F_{1}\cup F_{2}$ implies that $|F|\leq 4$ and that $H^{\prime}=H^{\prime}_{1}\cup H^{\prime}_{2}\cup F$ is spanning and connected in $G$. Moreover, observe that for $i\in\{1,2\}$ each edge $e\in H^{\prime}_{i}\cup F_{i}$ is either in a 2EC component of $H^{\prime}_{i}\cup F_{i}$ or lies on an $x-y$ path in $H^{\prime}_{i}\cup F_{i}$, where $x,y\in\{u,v,w\}$, $x\neq y$. Thus, Lemma 25 implies that $H^{\prime}$ is a 2EC spanning subgraph of $G$. ∎
Lemma 27.
If the condition in Algorithm 2 in Algorithm 2 applies and $\mathsf{RED}(G_{2}^{\mathsf{B1}})$ is a 2EC spanning subgraph of $G_{2}^{\mathsf{B1}}$, then there exist edges $F^{\mathsf{B1}}$ with $|F^{\mathsf{B1}}|\leq 1$ such that $H^{\mathsf{B1}}$ is a 2EC spanning subgraph of $G$.
Proof.
Assume w.l.o.g. that $u$ and $v$ belong to the same 2EC component of $\mathsf{OPT}_{1}^{\mathsf{B1}}$. Let $H_{1}^{\mathsf{B1}}\coloneq\mathsf{OPT}_{1}^{\mathsf{B1}}$ and $C_{i}^{\mathsf{B1}}(x)$ be the 2EC component of $H_{i}^{\mathsf{B1}}$ that contains $x$, where $x\in\{u,v,w\}$ and $i\in\{1,2\}$. Since $C_{1}^{\mathsf{B1}}(u)=C_{1}^{\mathsf{B1}}(v)$, we have that in $\mathsf{OPT}_{1}^{\mathsf{B1}}\cup H_{2}^{\mathsf{B1}}$, the vertices of $C_{1}^{\mathsf{B1}}(u)$, $C_{2}^{\mathsf{B1}}(u)$, and $C_{2}^{\mathsf{B1}}(v)$ are in the same 2EC component. Furthermore, in $\mathsf{OPT}_{1}^{\mathsf{B1}}$, there is a connection between $C_{1}^{\mathsf{B1}}(u)$ and $C_{1}^{\mathsf{B1}}(w)$. If $w$ and $v$ are already connected in $H_{2}^{\mathsf{B1}}$ then we can set $F^{\mathsf{B1}}\coloneq\emptyset$ :- ∅ and see that this satisfies the lemma, since $H^{\mathsf{B1}}$ contains a spanning tree and every edge in $H^{\mathsf{B1}}$ is either in the 2EC component containing $u$ and $v$, in a 2EC component containing $w$ or on a $w-v$ path in $G_{i}$, $i\in\{1,2\}$. Otherwise, if $w$ and $v$ are not in the same connected component in $H_{2}^{\mathsf{B1}}$, by Lemma 24, there must be an edge $e$ from $C_{2}^{\mathsf{B1}}(w)$ to either $C_{2}^{\mathsf{B1}}(v)$ or $C_{2}^{\mathsf{B1}}(u)$ in $E(G_{2})\setminus H^{\mathsf{B1}}$. Now it can be easily verified that $F^{\mathsf{B1}}\coloneq\{e\}$ satisfies the lemma, since $H^{\mathsf{B1}}$ contains a spanning tree and every edge in $H^{\mathsf{B1}}$ is either in the 2EC component containing $u$ and $v$, in a 2EC component containing $w$ or on a $w-v$ path in $G_{i}$, $i\in\{1,2\}$. ∎
Lemma 28.
If the condition in Algorithm 2 in Algorithm 2 applies and $\mathsf{RED}(G_{2}^{\mathsf{B2}})$ is a 2EC spanning subgraph of $G_{2}^{\mathsf{B2}}$, then there exist edges $F^{\mathsf{B2}}$ with $|F^{\mathsf{B2}}|\leq 2$ such that $H^{\mathsf{B2}}$ is a 2EC spanning subgraph of $G$.
Proof.
The proof is similar to the previous one. Assume w.l.o.g. that $u$ and $v$ belong to the same 2EC component of $\mathsf{OPT}_{1}^{\mathsf{B2}}$. Let $H_{1}^{\mathsf{B2}}\coloneq\mathsf{OPT}_{1}^{\mathsf{B2}}$ and $C_{i}^{\mathsf{B2}}(x)$ be the 2EC component of $H_{i}^{\mathsf{B2}}$ that contains $x$, where $x\in\{u,v,w\}$ and $i\in\{1,2\}$. Since $C_{1}^{\mathsf{B2}}(u)=C_{1}^{\mathsf{B2}}(v)$, we have that in $\mathsf{OPT}_{1}^{\mathsf{B2}}\cup H_{2}^{\mathsf{B2}}$, the vertices of $C_{1}^{\mathsf{B2}}(u)$, $C_{2}^{\mathsf{B2}}(u)$ and $C_{2}^{\mathsf{B2}}(v)$ are in the same 2EC component. Furthermore, in $\mathsf{OPT}_{1}^{\mathsf{B2}}$, there is no connection between $C_{1}^{\mathsf{B1}}(u)$ and $C_{1}^{\mathsf{B2}}(w)$, since it is of type $\mathsf{B2}$. Hence, by Lemma 24, there must be an edge $e_{1}\in E(G_{1})\setminus\mathsf{OPT}_{1}^{\mathsf{B2}}$. Note that $\mathsf{OPT}_{1}^{\mathsf{B2}}\cup\{e_{1}\}$ is now a solution of type $\mathsf{B1}$ for $G_{1}$, and hence we can do the same as in the proof of the previous lemma and observe that $\lvert F^{\mathsf{B2}}\rvert\leq 2$. ∎
Lemma 29.
If the condition in Algorithm 2 in Algorithm 2 applies and $\mathsf{RED}(G_{2}^{\mathsf{C1}})$ is a 2EC spanning subgraph of $G_{2}^{\mathsf{C1}}$, then there exist edges $F^{\mathsf{C1}}$ with $|F^{\mathsf{C1}}|\leq 2$ such that $H^{\mathsf{C1}}$ is a 2EC spanning subgraph of $G$.
Proof.
Let $C_{2}^{\mathsf{C1}}(x)$ be the 2EC component of $H_{2}^{\mathsf{C1}}$ that contains $x$, for $x\in\{u,v,w\}$. Note that $\mathsf{OPT}_{1}^{\mathsf{C1}}$ is connected and spanning for $G_{1}$ and that each edge $e\in\mathsf{OPT}_{1}^{\mathsf{C1}}$ is either in some 2EC component or it is on a $y-z$-path in $\mathsf{OPT}_{1}^{\mathsf{C1}}$ for $y,z\in\{u,v,w\}$, $y\neq z$. We consider three cases. If $H_{2}^{\mathsf{C1}}$ is connected, then we claim that $F^{\mathsf{C1}}\coloneq\emptyset$ :- ∅ satisfies the lemma. In this case every edge of $H_{2}^{\mathsf{C1}}$ is part of a 2EC component of $H_{2}^{\mathsf{C1}}$ or it is on a $y-z$-path in $H_{2}^{\mathsf{C1}}$ for $y,z\in\{u,v,w\}$, $y\neq z$. Hence, Lemma 25 implies that $H_{2}^{\mathsf{C1}}\cup\mathsf{OPT}_{1}^{\mathsf{C1}}$ is a 2EC spanning subgraph of $G$.
Next, assume that $H_{2}^{\mathsf{C1}}$ is composed of two connected components, say w.l.o.g. one containing $u$ and $v$ and the other containing $w$, and observe that each component is 2EC. Then Lemma 24 guarantees the existence of an edge $e$ in $G_{2}$ between $C_{2}^{\mathsf{C1}}(w)$ and $C_{2}^{\mathsf{C1}}(u)$ such that $H_{2}^{\mathsf{C1}}\cup\{e\}$ is connected. Thus, we can analogously to the previous case show that $\mathsf{OPT}_{1}^{\mathsf{C1}}\cup H_{2}^{\mathsf{C1}}\cup F^{\mathsf{C1}}$ with $F^{\mathsf{C1}}\coloneq\{e\}$ is a 2EC spanning subgraph of $G$, since in $H_{2}^{\mathsf{C1}}\cup F^{\mathsf{C1}}$ each edge is either in some 2EC component or it is on a $y-z$-path in $H_{2}^{\mathsf{C1}}\cup F^{\mathsf{C1}}$ for $y,z\in\{u,v,w\}$, $y\neq z$ (hence Lemma 25 holds). In the remaining case $H_{2}^{\mathsf{C1}}$ is composed of three connected components, each containing exactly one vertex of $\{u,v,w\}$. Observe that each component is 2EC. Hence, Lemma 24 guarantees the existence of at least two edges $e_{1}$ and $e_{2}$ in $G_{2}$ between two pairs of connected components of $H_{2}^{\mathsf{C1}}$ such that $H_{2}^{\mathsf{C1}}\cup\{e_{1},e_{2}\}$ is connected. Thus, we can analogously to the previous case(s) show that $\mathsf{OPT}_{1}^{\mathsf{C1}}\cup H_{2}^{\mathsf{C1}}\cup F^{\mathsf{C1}}$ with $F^{\mathsf{C1}}\coloneq\{e_{1},e_{2}\}$ is a 2EC spanning subgraph of $G$, since in $H_{2}^{\mathsf{C1}}\cup F^{\mathsf{C1}}$ each edge is either in some 2EC component or it is on a $y-z$-path in $H_{2}^{\mathsf{C1}}\cup F^{\mathsf{C1}}$ for $y,z\in\{u,v,w\}$, $y\neq z$ (hence Lemma 25 holds). ∎
Lemma 30.
If the condition in Algorithm 2 in Algorithm 2 applies and $\mathsf{RED}(G_{2}^{\mathsf{C2}})$ is a 2EC spanning subgraph of $G_{2}^{\mathsf{C2}}$, then there exist edges $F^{\mathsf{C2}}$ with $|F^{\mathsf{C2}}|\leq 1$ such that $H^{\mathsf{C2}}$ is a 2EC spanning subgraph of $G$ and $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|\leq-2$.
Proof.
Fix an optimal $\mathsf{C2}$ solution $\mathsf{OPT}_{1}^{\mathsf{C2}}$ for $G_{1}$. Note that both dummy edges $uy$ and $vy$ must be part of $\mathsf{RED}(G_{2}^{\mathsf{C2}})$ as otherwise $\mathsf{RED}(G_{2}^{\mathsf{C2}})$ cannot be feasible for $G_{2}^{\mathsf{C2}}$. By the same reason, $H_{2}^{\mathsf{C2}}$ must be connected and spanning for $G_{2}$.
We first consider the case where $vw\in\mathsf{RED}(G_{2}^{\mathsf{C2}})$. Let $C_{1}^{\mathsf{C2}}(w)$ be the 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ that contains $w$. Since $C_{1}^{\mathsf{C2}}(w)$ is isolated in $\mathsf{OPT}_{1}^{\mathsf{C2}}$, Lemma 24 implies that there exists an edge $e$ in $G_{1}$ between $C_{1}^{\mathsf{C2}}(w)$ and $V(G_{1})\setminus C_{1}^{\mathsf{C2}}(w)$. Then, $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ is connected and spanning for $G_{1}$, and every edge in $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ between any two distinct vertices of $\{u,v,w\}$ between which is a path in $H_{2}^{\mathsf{C2}}$. Similarly, every edge in $H_{2}^{\mathsf{C2}}$ is part of a 2EC component of $H_{2}^{\mathsf{C2}}$ or lies on a path in $H_{2}^{\mathsf{C2}}$ between any two distinct vertices of $\{u,v,w\}$ between which is a path in $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$. Thus, $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup H_{2}^{\mathsf{C2}}\cup F^{\mathsf{C2}}$ with $F^{\mathsf{C2}}\coloneq\{e\}$ is a 2EC spanning subgraph of $G$ by Lemma 25. Moreover, $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|=1-3=-2$.
If $vw\notin\mathsf{RED}(G_{2}^{\mathsf{C2}})$, then every edge in $\mathsf{OPT}_{1}^{\mathsf{C2}}$ is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C2}}$ between $u$ and $v$ between which is also a path in $H_{2}^{\mathsf{C2}}$. Further, every edge in $H_{2}^{\mathsf{C2}}$ is part of a 2EC component of $H_{2}^{\mathsf{C2}}$ or lies on a path in $H_{2}^{\mathsf{C2}}$ between $u$ and $v$, and there is a path between $u$ and $v$ in $\mathsf{OPT}_{1}^{\mathsf{C2}}$. Note that there cannot be an edge in $H_{2}^{\mathsf{C2}}$ that is on a path between $w$ and $u$ or $v$ but not in a 2EC component, because then $\mathsf{RED}(G_{2}^{\mathsf{C2}})$ cannot be feasible for $G_{2}^{\mathsf{C2}}$ as $vw\notin\mathsf{RED}(G_{2}^{\mathsf{C2}})$. Thus, $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup H_{2}^{\mathsf{C2}}$ is a 2EC spanning subgraph of $G$ by Lemma 25. We have that $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|=0-2=-2$. ∎
Lemma 31.
If the condition in Algorithm 2 in Algorithm 2 applies and $\mathsf{RED}(G_{2}^{\mathsf{C2}})$ is a 2EC spanning subgraph of $G_{2}^{\mathsf{C2}}$, then there exist edges $F^{\mathsf{C2}}$ with $|F^{\mathsf{C2}}|\leq 1$ and $\mathsf{OPT}_{1}^{\mathsf{C2}}$ such that $H^{\mathsf{C2}}$ is a 2EC spanning subgraph of $G$ and $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|\leq-3$.
Proof.
Let $D=\{uy,vz,zy,wy\}$. First observe that $H_{2}^{\mathsf{C2}}$ must be connected and spanning for $G_{2}$. Note that $vz$ and $zy$ must be part of $\mathsf{RED}(G_{2}^{\mathsf{C2}})$ to make $z$ incident to 2 edges, and additionally $vy$ or $wy$ must be part of $\mathsf{RED}(G_{2}^{\mathsf{C2}})$ to make $y$ incident to at least 2 edges. Thus, we can distinguish the following cases.
If $\mathsf{RED}(G_{2}^{\mathsf{C2}})\cap D=\{uy,vz,zy\}$ (or by symmetry $\mathsf{RED}(G_{2}^{\mathsf{C2}})\cap D=\{wy,vz,zy\}$), let $\mathsf{OPT}_{1}^{\mathsf{C2}}$ be an optimal $\mathsf{C2}$ solution for $G_{1}$ that contains a $C_{1}^{\mathsf{C2}}(u)-C_{1}^{\mathsf{C2}}(v)$ path, where $C_{1}^{\mathsf{C2}}(x)$ denotes the 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ that contains $x$, for $x\in\{u,v,w\}$. Note that every edge of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C2}}$ between $u$ and $v$ between which is also a path in $H_{2}^{\mathsf{C2}}$. Further, every edge in $H_{2}^{\mathsf{C2}}$ is part of a 2EC component of $H_{2}^{\mathsf{C2}}$ or lies on a path in $H_{2}^{\mathsf{C2}}$ between $u$ and $v$, and there is a path between $u$ and $v$ in $\mathsf{OPT}_{1}^{\mathsf{C2}}$. Note that there cannot be an edge in $H_{2}^{\mathsf{C2}}$ that is on a path between $w$ and $u$ or $v$ but not in a 2EC component, because then $\mathsf{RED}(G_{2}^{\mathsf{C2}})$ cannot be feasible for $G_{2}^{\mathsf{C2}}$ as $wy\notin\mathsf{RED}(G_{2}^{\mathsf{C2}})$. Thus, $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup H_{2}^{\mathsf{C2}}$ is a 2EC spanning subgraph of $G$ by Lemma 25. We have that $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|=0-3=-3$.
Otherwise, that is, $\mathsf{RED}(G_{2}^{\mathsf{C2}})\cap D=\{vy,vz,zy,wy\}$, let $\mathsf{OPT}_{1}^{\mathsf{C2}}$ be an optimal $\mathsf{C2}$ solution for $G_{1}$ that contains a $C_{1}^{\mathsf{C2}}(u)-C_{1}^{\mathsf{C2}}(v)$ path of $\mathsf{OPT}_{1}^{\mathsf{C2}}$, where $C_{1}^{\mathsf{C2}}(x)$ denotes the 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ that contains $x$, for $x\in\{u,v,w\}$. Lemma 24 implies that there exists an edge $e$ in $G_{1}$ between $C_{1}^{\mathsf{C2}}(w)$ and the connected component containing $V(G_{1})\setminus V(C_{1}^{\mathsf{C2}}(w))$. Therefore $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ is connected. Thus, every edge in $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ (resp. $H_{2}^{\mathsf{C2}}$) is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ ($H_{2}^{\mathsf{C2}}$) or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ ($H_{2}^{\mathsf{C2}}$) between any two distinct vertices of $\{u,v,w\}$ between which is a path in $H_{2}^{\mathsf{C2}}$ ($\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$). We conclude using Lemma 25 that $H_{2}^{\mathsf{C2}}\cup\mathsf{OPT}_{1}^{\mathsf{C2}}\cup F^{\mathsf{C2}}$ with $F^{\mathsf{C2}}\coloneq\{e\}$ is a 2EC spanning subgraph of $G$. We have that $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|=1-4=-3$. ∎
Lemma 32.
If the condition in Algorithm 2 in Algorithm 2 applies and $\mathsf{RED}(G_{2}^{\mathsf{C2}})$ is a 2EC spanning subgraph of $G_{2}^{\mathsf{C2}}$, then there exist edges $F^{\mathsf{C2}}$ with $|F^{\mathsf{C2}}|\leq 1$ and $\mathsf{OPT}_{1}^{\mathsf{C2}}$ such that $H^{\mathsf{C2}}$ is a 2EC spanning subgraph of $G$ and $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|\leq-2$.
Proof.
Let $D=\{uy,vy,wy\}$. First observe that $H_{2}^{\mathsf{C2}}$ must be connected and spanning for $G_{2}$. Note that $|D\cap\mathsf{RED}(G_{2}^{\mathsf{C2}})|\geq 2$ as $y$ needs to be 2EC in $\mathsf{RED}(G_{2}^{\mathsf{C2}})$.
If $|D\cap\mathsf{RED}(G_{2}^{\mathsf{C2}})|=2$, let w.l.o.g. by symmetry $D\cap\mathsf{RED}(G_{2}^{\mathsf{C2}})=\{uy,vy\}$ and $\mathsf{OPT}_{1}^{\mathsf{C2}}$ be an optimal $\mathsf{C2}$ solution for $G_{1}$ that contains a $C_{1}^{\mathsf{C2}}(u)-C_{1}^{\mathsf{C2}}(v)$ path, where $C_{i}^{\mathsf{C2}}(x)$ denotes the 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ that contains $x$, for $x\in\{u,v,w\}$. Note that every edge of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C2}}$ between $u$ and $v$ between which is also a path in $H_{2}^{\mathsf{C2}}$. Further, every edge in $H_{2}^{\mathsf{C2}}$ is part of a 2EC component of $H_{2}^{\mathsf{C2}}$ or lies on a path in $H_{2}^{\mathsf{C2}}$ between $u$ and $v$, and there is a path between $u$ and $v$ in $\mathsf{OPT}_{1}^{\mathsf{C2}}$. Note that there cannot be an edge in $H_{2}^{\mathsf{C2}}$ that is on a path between $w$ and $u$ or $v$ but not in a 2EC component, because then $\mathsf{RED}(G_{2}^{\mathsf{C2}})$ cannot be feasible for $G_{2}^{\mathsf{C2}}$ as $wy\notin\mathsf{RED}(G_{2}^{\mathsf{C2}})$. Thus, $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup H_{2}^{\mathsf{C2}}$ is a 2EC spanning subgraph of $G$ by Lemma 25. We have that $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|=0-2=-2$.
Otherwise, that is, $\mathsf{RED}(G_{2}^{\mathsf{C2}})\cap D=\{uy,vy,wy\}$, let $\mathsf{OPT}_{1}^{\mathsf{C2}}$ be an optimal $\mathsf{C2}$ solution for $G_{1}$ that contains a $C_{1}^{\mathsf{C2}}(u)-C_{1}^{\mathsf{C2}}(v)$ path, where $C_{i}^{\mathsf{C2}}(x)$ denotes the 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ that contains $x$, for $x\in\{u,v,w\}$. Lemma 24 implies that there exists an edge $e$ in $G_{1}$ between $C_{1}^{\mathsf{C2}}(w)$ and the connected component containing $V(G_{1})\setminus V(C_{1}^{\mathsf{C2}}(w))$. Therefore $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ is connected. Thus, every edge in $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ (resp. $H_{2}^{\mathsf{C2}}$) is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ ($H_{2}^{\mathsf{C2}}$) or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ ($H_{2}^{\mathsf{C2}}$) between any two distinct vertices of $\{u,v,w\}$ between which is a path in $H_{2}^{\mathsf{C2}}$ ($\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$). We conclude using Lemma 25 that $H_{2}^{\mathsf{C2}}\cup\mathsf{OPT}_{1}^{\mathsf{C2}}\cup F^{\mathsf{C2}}$ with $F^{\mathsf{C2}}\coloneq\{e\}$ is a 2EC spanning subgraph of $G$. We have that $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|=1-3=-2$. ∎
Lemma 33.
If the condition in Algorithm 2 in Algorithm 2 applies and $\mathsf{RED}(G_{2}^{\mathsf{C3}})$ is a 2EC spanning subgraph of $G_{2}^{\mathsf{C3}}$, then there exist edges $F^{\mathsf{C3}}$ with $|F^{\mathsf{C3}}|\leq 4$ such that $H^{\mathsf{C3}}$ is a 2EC spanning subgraph of $G$ and $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|\leq 0$.
Proof.
First note that the claimed $\mathsf{C3}$ solution in Algorithm 2 is guaranteed by Lemma 24 (among renaming $u$, $v$ and $w$). Let $C_{2}^{\mathsf{C3}}(x)$ be the 2EC component of $H_{2}^{\mathsf{C3}}$ that contains $x$, for $x\in\{u,v,w\}$. We have 4 main cases.
Case 1: $|\mathsf{RED}(G_{2}^{\mathsf{C3}})\cap\{uv,vw\}|=\ell\leq 2$. First observe that $H_{2}^{\mathsf{C3}}$ is connected as otherwise $\mathsf{RED}(G_{2}^{\mathsf{C3}})$ cannot be feasible for $G_{2}^{\mathsf{C3}}$. Moreover, we can select $\ell$ edges $F^{\mathsf{C3}}\subseteq\{e^{\mathsf{C3}}_{uv},e^{\mathsf{C3}}_{vw}\}$ such that $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ is connected. Further note that every edge in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ (resp. $H_{2}^{\mathsf{C3}}$) is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ ($H_{2}^{\mathsf{C3}}$) or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ ($H_{2}^{\mathsf{C3}}$) between any two distinct vertices of $\{u,v,w\}$ between which is a path in $H_{2}^{\mathsf{C3}}$ ($\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$). Thus, Lemma 25 implies that $H_{2}^{\mathsf{C3}}\cup\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ is a 2EC spanning subgraph of $G$. We have that $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|=\ell-\ell=0$.
Case 2: $|\mathsf{RED}(G_{2}^{\mathsf{C3}})\cap\{uv,uv\}|=2$ (the case $|\mathsf{RED}(G_{2}^{\mathsf{C3}})\cap\{vw,vw\}|=2$ is symmetric). We have that $C_{2}^{\mathsf{C3}}(v)=C_{2}^{\mathsf{C3}}(w)$, as otherwise $\mathsf{RED}(G_{2}^{\mathsf{C3}})$ cannot be feasible for $G_{2}^{\mathsf{C3}}$. If $H_{2}^{\mathsf{C3}}$ is 2EC, then clearly $H_{2}^{\mathsf{C3}}\cup\mathsf{OPT}_{2}^{\mathsf{C3}}$ is a 2EC spanning subgraph of $G$, and $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|=0-2\leq 0$.
If $C_{2}^{\mathsf{C3}}(u)$ is connected to $C_{2}^{\mathsf{C3}}(v)$ in $H_{2}^{\mathsf{C3}}$, then $H_{2}^{\mathsf{C3}}$ is connected. Further, $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ with $F^{\mathsf{C3}}\coloneq\{e^{\mathsf{C3}}_{uv}\}$ is connected. Every edge in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ between $u$ and $v$, which are also connected in $H_{2}^{\mathsf{C3}}$. Note that there cannot be an edge in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ on a path between $w$ and $u$ or $v$ that is not in a 2EC component by the choice of the $\mathsf{C3}$ solution $\mathsf{OPT}_{1}^{\mathsf{C3}}$. Moreover, every edge in $H_{2}^{\mathsf{C3}}$ is part of a 2EC component of $H_{2}^{\mathsf{C3}}$ or lies on a path in $H_{2}^{\mathsf{C3}}$ between $u$ and $v$ or $w$, which are also connected in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$. Note that there cannot be an edge in $H_{2}^{\mathsf{C3}}$ on a path between $v$ and $w$ that is not in a 2EC component, because $C_{2}^{\mathsf{C3}}(v)=C_{2}^{\mathsf{C3}}(w)$ is 2EC. Thus, Lemma 25 implies that $H_{2}^{\mathsf{C3}}\cup\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ is a 2EC spanning subgraph of $G$, and $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|=1-2\leq 0$.
Otherwise, that is, $C_{2}^{\mathsf{C3}}(u)$ is not connected to $C_{2}^{\mathsf{C3}}(v)$ in $H_{2}^{\mathsf{C3}}$, by Lemma 24 there must exist an edge $e$ in $G_{2}$ between $C_{2}^{\mathsf{C3}}(u)$ and $C_{2}^{\mathsf{C3}}(v)$ such that $H_{2}^{\mathsf{C3}}\cup\{e\}$ is connected. Thus, we can analogously to the previous case show that $H_{2}^{\mathsf{C3}}\cup\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ with $F^{\mathsf{C3}}\coloneq\{e,e^{\mathsf{C3}}_{uv}\}$ is a 2EC spanning subgraph of $G$. We have that $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|=2-2=0$.
Case 3: $|\mathsf{RED}(G_{2}^{\mathsf{C3}})\cap\{uv,uv,vw\}|=3$ (the case $|\mathsf{RED}(G_{2}^{\mathsf{C3}})\cap\{uv,vw,vw\}|=3$ is symmetric). Note that in this case $C_{2}^{\mathsf{C3}}(u)$ is connected to $C_{2}^{\mathsf{C3}}(v)$ in $H_{2}^{\mathsf{C3}}$ as otherwise $u$ and $w$ cannot be 2EC in $\mathsf{RED}(G_{2}^{\mathsf{C3}})$. We have two cases.
If $C_{2}^{\mathsf{C3}}(v)$ is connected to $C_{2}^{\mathsf{C3}}(w)$ in $H_{2}^{\mathsf{C3}}$, then $H_{2}^{\mathsf{C3}}$ is connected. Further, $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ with $F^{\mathsf{C3}}\coloneq\{e^{\mathsf{C3}}_{uv},e^{\mathsf{C3}}_{vw}\}$ is connected and spanning for $G_{1}$. Thus, every edge in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ (resp. $H_{2}^{\mathsf{C3}}$) is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ ($H_{2}^{\mathsf{C3}}$) or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ ($H_{2}^{\mathsf{C3}}$) between any two distinct vertices of $\{u,v,w\}$ between which is a path in $H_{2}^{\mathsf{C3}}$ ($\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$). We conclude via Lemma 25 that $H_{2}^{\mathsf{C3}}\cup\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ is a 2EC spanning subgraph of $G$, and $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|=2-3\leq 0$.
Otherwise, that is, $C_{2}^{\mathsf{C3}}(v)$ is not connected to $C_{2}^{\mathsf{C3}}(w)$ in $H_{2}^{\mathsf{C3}}$, Lemma 24 guarantees the existence of an edge $e$ in $G_{2}$ between the vertices of $V(C_{2}^{\mathsf{C3}}(w))$ and the vertices of $V(H_{2}^{\mathsf{C3}}\setminus V(C_{2}^{\mathsf{C3}}(w))$. Therefore, $H_{2}^{\mathsf{C3}}\cup\{e\}$ is connected. Hence, we can analogously to the previous case show that $H_{2}^{\mathsf{C3}}\cup\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ with $F^{\mathsf{C3}}\coloneq\{e,e^{\mathsf{C3}}_{uv},e^{\mathsf{C3}}_{vw}\}$ is a 2EC spanning subgraph of $G$. We have that $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|=3-3=0$.
Case 4: $|\mathsf{RED}(G_{2}^{\mathsf{C3}})\cap\{uv,uv,vw,vw\}|=4$. We distinguish three cases.
First, assume that $H_{2}^{\mathsf{C3}}$ is connected. We have that $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ with $F^{\mathsf{C3}}\coloneq\{e^{\mathsf{C3}}_{uv},e^{\mathsf{C3}}_{vw}\}$ is connected and spanning for $G_{1}$. Thus, every edge in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ (resp. $H_{2}^{\mathsf{C3}}$) is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ ($H_{2}^{\mathsf{C3}}$) or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ ($H_{2}^{\mathsf{C3}}$) between any two distinct vertices of $\{u,v,w\}$ between which is a path in $H_{2}^{\mathsf{C3}}$ ($\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$). We conclude via Lemma 25 that $H_{2}^{\mathsf{C3}}\cup\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ is a 2EC spanning subgraph of $G$, and $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|=2-4\leq 0$.
If $H_{2}^{\mathsf{C3}}$ is composed of two connected components, assume w.l.o.g. that $C_{2}^{\mathsf{C3}}(u)$ is connected to $C_{2}^{\mathsf{C3}}(v)$ in $H_{2}^{\mathsf{C3}}$ and the other connected component in $H_{2}^{\mathsf{C3}}$ is $C_{2}^{\mathsf{C3}}(w)$. In this case, Lemma 24 guarantees that there is an edge $e$ in $E(G_{2})\setminus C_{2}^{\mathsf{C3}}(w)$ between the vertices of $C_{2}^{\mathsf{C3}}(w)$ and the vertices of $V(H_{2}^{\mathsf{C3}})\setminus V(C_{2}^{\mathsf{C3}}(w))$. Therefore, $H_{2}^{\mathsf{C3}}\cup\{e\}$ is connected. Hence, we can analogously to the previous case show that $H_{2}^{\mathsf{C3}}\cup\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ with $F^{\mathsf{C3}}\coloneq\{e,e^{\mathsf{C3}}_{uv},e^{\mathsf{C3}}_{vw}\}$ is a 2EC spanning subgraph of $G$, and $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|=3-4\leq 0$.
Otherwise, that is, $C_{2}^{\mathsf{C3}}(u)$, $C_{2}^{\mathsf{C3}}(v)$ and $C_{2}^{\mathsf{C3}}(w)$ are all pairwise disjoint, Lemma 24 guarantees the existence of two edges $F\subseteq E(G_{2})$ such that $H_{2}^{\mathsf{C3}}\cup F$ is connected. Hence, we can analogously to the previous case(s) show that $H_{2}^{\mathsf{C3}}\cup\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ with $F^{\mathsf{C3}}\coloneq F\cup\{e^{\mathsf{C3}}_{uv},e^{\mathsf{C3}}_{vw}\}$ is a 2EC spanning subgraph of $G$. We have that $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|=4-4=0$. ∎
A.3 Final Lemmas
In this final subsection, we prove Definition 5. The statement is implied by the following three lemmas. We first show that $\mathsf{RED}(G)$ runs in polynomial time. The proof is similar to the corresponding proof in [18].
Lemma 34 (running time).
For any 2-ECSS instance $G$, $\mathsf{RED}(G)$ runs in polynomial time in $|V(G)|$ if $\mathsf{ALG}(G)$ runs in polynomial time in $|V(G)|$.
Proof.
Let $n=|V(G)|$ and $m=|V(G)|$, and $\sigma=n^{2}+m^{2}$ be the size of the problem. We first argue that every non-recursive step of Algorithms 1 and 2 can be performed in time polynomial in $\sigma$. For all computations performed in Algorithm 1 this follows from [18]. We already argued that the computation in Algorithm 2 in Algorithm 2 can be executed in polynomial time in $\sigma$. Moreover, determining the $\mathsf{C2}$ subcases in Algorithms 2, 2 and 2 and selecting a desired $\mathsf{OPT}_{1}^{\mathsf{C2}}$ solution in Algorithms 2 and 2 and $\mathsf{OPT}_{1}^{\mathsf{C3}}$ solution in Algorithm 2 can be done in constant time, as $G_{1}$ is of constant size in this case. Finding the corresponding sets of edges $F$ in Algorithm 2 can also be done in polynomial time in $\sigma$ by enumeration, because Lemmas 26, 27, 28, 29, 30, 31, 32 and 33 guarantee that there exists a desired set of edges $F$ of constant size.
To bound the total running time including recursive steps, let $T(\sigma)$ be the running time of Algorithm 1 as a function of $\sigma$. In each step, starting with a problem of size $\sigma$, we generate either a single subproblem of size $\sigma^{\prime}<\sigma$ (to easily see this in Algorithm 2 recall that $7\leq|V_{1}|\leq|V_{2}|$) or two subproblems of size $\sigma^{\prime}$ and $\sigma^{\prime\prime}$ such that $\sigma^{\prime},\sigma^{\prime\prime}<\sigma$ and $\sigma^{\prime}+\sigma^{\prime\prime}\leq\sigma$. Thus, $T(\sigma)\leq\max\{T(\sigma^{\prime}),T(\sigma^{\prime})+T(\sigma^{\prime\prime})\}+\mathrm{poly}(\sigma)$ with $\sigma^{\prime},\sigma^{\prime\prime}<\sigma$ and $\sigma^{\prime}+\sigma^{\prime\prime}\leq\sigma$. Therefore, $T(\sigma)\leq\sigma\cdot\mathrm{poly}(\sigma)$. ∎
Next, we show that $\mathsf{RED}(G)$ returns a feasible solution to the 2-ECSS problem on $G$.
Lemma 35 (correctness).
For any 2-ECSS instance $G$, $\mathsf{RED}(G)$ returns a 2EC spanning subgraph of $G$.
Proof.
The proof is by induction over the tuple $(\lvert V(G)\rvert,\lvert E(G)\rvert)$. The base cases are given by Algorithms 1 and 1 of Algorithm 1. If we enter Algorithm 1, the claim trivially holds. If we call $\mathsf{ALG}$ in Algorithm 1, the claim follows by the correctness of $\mathsf{ALG}$ for $(\alpha,\varepsilon)$-structured instances. If in Algorithm 1 any of the conditions Algorithms 1, 1, 1, 1 and 1 apply, the claim follows from [18, Lemma A.6].
Thus, for the remaining proof we consider the case where Algorithm 1 holds and we enter Algorithm 2. In this case the correctness follows from the induction hypothesis, Corollary 23 and Lemmas 27, 28, 29, 30, 31, 32 and 33. ∎
Let $\mathsf{red}(G)\coloneq|\mathsf{RED}(G)|$ denote the size of the solution computes by $\mathsf{RED}$ for a given 2-ECSS instance $G$. The following lemma implies $\mathsf{red}(G)\leq(\alpha+4\varepsilon)\mathsf{opt}(G)$ using the trivial bound $\mathsf{opt}(G)\geq|V(G)|$. While its proof is quite long, it is not difficult and easy to verify: We consider different cases depending on the reduction the algorithm uses and, if Algorithm 2 is executed, we distinguish for some cases further which solution types the actual optimal solution uses for both sides of the cut. For each case, we can easily derive bounds for the different parts of the solution that our algorithm produces, and put them together to obtain the overall bound, which is stated below.
Lemma 36 (approximation guarantee).
For any 2-ECSS instance $G$ it holds that
$\mathsf{red}(G)\leq\begin{cases}\mathsf{opt}(G)&\quad\text{if }\lvert V(G)\rvert\leq\frac{4}{\varepsilon}\ ,\text{ and}\\ \alpha\cdot\mathsf{opt}(G)+4\varepsilon\cdot\lvert V(G)\rvert-4&\quad\text{if }\lvert V(G)\rvert>\frac{4}{\varepsilon}\ .\end{cases}$
Proof.
The proof is by induction over the tuple $(\lvert V(G)\rvert,\lvert E(G)\rvert)$. The base cases are given by Algorithms 1 and 1 of Algorithm 1. If we enter Algorithm 1, the stated bound trivially holds. If we call $\mathsf{ALG}$ in Algorithm 1, the claim follows because we assume that $\mathsf{ALG}(G)$ returns an $\alpha$-approximate solution to the 2-ECSS problem on $G$ and $G$ is in this case $(\alpha,\varepsilon)$-structured by the correctness of $\mathsf{RED}$ (cf. Lemma 35).
If in Algorithm 1 any of the reductions in Algorithms 1, 1, 1, 1 and 1 is executed, the claim follows from [18, Lemma A.7]. (The proof in [18] proves the reduction with even stronger assumptions, which can be relaxed to our constants.)
If Algorithm 2 is executed, let $\mathsf{OPT}_{i}\coloneq\mathsf{OPT}(G)\cap E(G_{i})$ and $\mathsf{opt}_{i}\coloneq\lvert\mathsf{OPT}_{i}\rvert$ for $i\in\{1,2\}$. Note that $\mathsf{opt}(G)=\mathsf{opt}_{1}+\mathsf{opt}_{2}$. For the remainder of this proof, we only consider the execution of Algorithm 2. In particular, all line references (if not specified differently) refer to Algorithm 2. First, suppose that the condition in Algorithm 2 holds. Note that for $i\in\{1,2\}$ we have that $\mathsf{OPT}_{i}$ is a feasible solution for $G_{i}^{\prime}$ and therefore $\mathsf{opt}_{i}\geq\mathsf{opt}(G_{i}^{\prime})$. Using the induction hypothesis, Lemma 26 and $\lvert V(G_{1})\rvert+\lvert V(G_{2})\rvert=\lvert V_{1}\rvert+\lvert V_{2}\rvert+2=\lvert V(G)\rvert-1$, we conclude that
$\mathsf{red}(G)=\lvert H^{\prime}_{1}\rvert+\lvert H^{\prime}_{2}\rvert+\lvert F^{\prime}\rvert\leq\alpha\cdot(\mathsf{opt}_{1}+\mathsf{opt}_{2})+4\varepsilon(\lvert V_{1}\rvert+\lvert V_{2}\rvert+2)-8+4\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\cdot(\lvert V(G)\rvert)-4\ .$
It remains to verify the case where we reach Algorithm 2. We distinguish several cases depending on which case the algorithm uses. Note that Lemma 21 gives $\mathsf{opt}_{1}\geq\mathsf{opt}_{1}^{\min}\geq 8$. Furthermore, throughout the proof we require $\alpha\geq\frac{5}{4}$. We use both statements in every case below.
(Algorithm 2: $\mathsf{OPT}_{1}^{\mathsf{B1}}$ exists and $\mathsf{opt}_{1}^{\mathsf{B1}}\leq\mathsf{opt}_{1}^{\min}+1$)
Note that $\mathsf{opt}(G_{2}^{\mathsf{B1}})\leq\mathsf{opt}_{2}$, because $u$, $v$, and $w$ are contracted into a single vertex in $G_{2}^{\mathsf{B1}}$. Thus, $\mathsf{opt}(G)=\mathsf{opt}_{2}+\mathsf{opt}_{1}\geq\mathsf{opt}(G_{2}^{\mathsf{B1}})+\mathsf{opt}_{1}$. Moreover, $\mathsf{opt}_{1}^{\mathsf{B1}}\leq\mathsf{opt}_{1}^{\min}+1\leq\mathsf{opt}_{1}+1$, and $\lvert F^{\mathsf{B1}}\rvert\leq 1$ by Lemma 27. Therefore, we conclude
$\mathsf{red}(G)=\lvert H^{\mathsf{B1}}\rvert$ $\leq\mathsf{opt}_{1}^{\mathsf{B1}}+\lvert H_{2}^{\mathsf{B1}}\rvert+\lvert F^{\mathsf{B1}}\rvert\leq(\mathsf{opt}_{1}+1)+(\alpha\cdot\mathsf{opt}_{2}(G_{2}^{\mathsf{B1}})+4\varepsilon\lvert V(G_{2}^{\mathsf{B1}})\rvert-4)+1$
$\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4-(\alpha-1)\mathsf{opt}_{1}+2\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
(Algorithms 2 and 2: $t^{\min}=\mathsf{B2}$ or $t^{\min}=\mathsf{C1}$)
Let $\tau\coloneq t^{\min}$. Note that $\mathsf{opt}(G_{2}^{\tau})\leq\mathsf{opt}_{2}$, because $u$, $v$, and $w$ are contracted into a single vertex in $G_{2}^{\tau}$. Thus, $\mathsf{opt}(G)\geq\mathsf{opt}_{2}+\mathsf{opt}_{1}(G_{1}^{\tau})\geq\mathsf{opt}(G_{2}^{\tau})+\mathsf{opt}_{1}(G_{1}^{\tau})$. Moreover, $\lvert F^{\tau}\rvert\leq 2$ by Lemmas 28 and 29. Therefore, we conclude
$\mathsf{red}(G)=\lvert H^{\tau}\rvert$ $\leq\mathsf{opt}_{1}^{\tau}+\lvert H_{2}^{\tau}\rvert+\lvert F^{\tau}\rvert\leq\mathsf{opt}^{\tau}_{1}+(\alpha\cdot\mathsf{opt}_{2}(G_{2}^{\tau})+4\varepsilon\lvert V(G_{2}^{\tau})\rvert-4)+2$
$\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4-(\alpha-1)\mathsf{opt}^{\tau}_{1}+2\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
(Algorithm 2: $t^{\min}=\mathsf{C2}$)
In Algorithm 2 we have the following $\mathsf{C2}$ subtypes, which use different reductions $G_{2}^{\mathsf{C2}}$ for the recursion (Algorithms 2, 2 and 2).
Subtype $\mathsf{C2(i)}$ (Algorithm 2):
Every $\mathsf{OPT}_{1}^{\mathsf{C2}}$ solution contains a $C(u)-C(v)$ path. In this case the definition of Algorithm 2, Lemma 30, and the induction hypothesis gives
$\mathsf{red}(G)=|H^{\mathsf{C2}}|$ $\leq|\mathsf{OPT}_{1}^{\mathsf{C2}}|+\lvert H_{2}^{\mathsf{C2}}\rvert+\lvert F^{\mathsf{C2}}\rvert\leq\mathsf{opt}_{1}^{\mathsf{C2}}+\mathsf{red}(G_{2}^{\mathsf{C2}})-2$
$\leq\mathsf{opt}_{1}^{\mathsf{C2}}+\alpha\cdot\mathsf{opt}(G_{2}^{\mathsf{C2}})+4\varepsilon\lvert V(G_{2}^{\mathsf{C2}})\rvert-4-2\ .$ (1)
Subtype $\mathsf{C2(ii)}$ (Algorithm 2):
Every $\mathsf{OPT}_{1}^{\mathsf{C2}}$ solution contains either a $C(u)-C(v)$ path or a $C(v)-C(w)$ path. In this case the definition of Algorithm 2, Lemma 31, and the induction hypothesis gives
$\mathsf{red}(G)=|H^{\mathsf{C2}}|$ $\leq|\mathsf{OPT}_{1}^{\mathsf{C2}}|+\lvert H_{2}^{\mathsf{C2}}\rvert+\lvert F^{\mathsf{C2}}\rvert\leq\mathsf{opt}_{1}^{\mathsf{C2}}+\mathsf{red}(G_{2}^{\mathsf{C2}})-3$
$\leq\mathsf{opt}_{1}^{\mathsf{C2}}+\alpha\cdot\mathsf{opt}(G_{2}^{\mathsf{C2}})+4\varepsilon\lvert V(G_{2}^{\mathsf{C2}})\rvert-4-3\ .$ (2)
Subtype $\mathsf{C2(iii)}$ (Algorithm 2):
Every $\mathsf{OPT}_{1}^{\mathsf{C2}}$ solution contains some path between $C(u)$, $C(v)$, and $C(w)$. In this case the definition of Algorithm 2, Lemma 32, and the induction hypothesis gives
$\mathsf{red}(G)=|H^{\mathsf{C2}}|$ $\leq|\mathsf{OPT}_{1}^{\mathsf{C2}}|+\lvert H_{2}^{\mathsf{C2}}\rvert+\lvert F^{\mathsf{C2}}\rvert\leq\mathsf{opt}_{1}^{\mathsf{C2}}+\mathsf{red}(G_{2}^{\mathsf{C2}})-2$
$\leq\mathsf{opt}_{1}^{\mathsf{C2}}+\alpha\cdot\mathsf{opt}(G_{2}^{\mathsf{C2}})+4\varepsilon\lvert V(G_{2}^{\mathsf{C2}})\rvert-4-2\ .$ (3)
We distinguish the following cases depending on which type combination of Proposition 18 is present for $(\mathsf{OPT}_{1},\mathsf{OPT}_{2})$.
  1. (a)
    $(\mathsf{A},\{\mathsf{A},\mathsf{B1},\mathsf{B2},\mathsf{C1},\mathsf{C2},\mathsf{C3}\})$. Since $\mathsf{OPT}_{1}^{\mathsf{A}}$ exists, Lemma 22 gives $\mathsf{opt}_{1}=\mathsf{opt}_{1}^{\mathsf{A}}\geq\mathsf{opt}_{1}^{\mathsf{C2}}+3$.
    Subtype $\mathsf{C2(i)}$:
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+5$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vy,vw\}$ and at most two more edges, which are guaranteed by Lemma 24 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, (1) gives
    $\mathsf{red}(G)$ $\leq(\mathsf{opt}_{1}-3)+(\alpha\cdot(\mathsf{opt}_{2}+5)+4\varepsilon\lvert V(G_{2}^{\mathsf{C2}})\rvert-4-2$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(5-\mathsf{opt}_{1})\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
    Subtype $\mathsf{C2(ii)}$:
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+6$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vz,zy,wy\}$ and at most two more edges, which are guaranteed by Lemma 24 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, (2) gives
    $\mathsf{red}(G)$ $\leq(\mathsf{opt}_{1}-3)+(\alpha\cdot(\mathsf{opt}_{2}+6)+4\varepsilon\lvert V(G_{2}^{\mathsf{C2}})\rvert-4-3$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(6-\mathsf{opt}_{1})\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
    Subtype $\mathsf{C2(iii)}$:
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+5$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vy,wy\}$ and at most two more edges, which are guaranteed by Lemma 24 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, (3) gives
    $\mathsf{red}(G)$ $\leq(\mathsf{opt}_{1}-3)+(\alpha\cdot(\mathsf{opt}_{2}+5)+4\varepsilon\lvert V(G_{2}^{\mathsf{C2}})\rvert-4-2$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(5-\mathsf{opt}_{1})\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
  2. (b)
    $(\mathsf{B1},\{\mathsf{A},\mathsf{B1},\mathsf{B2},\mathsf{C1},\mathsf{C2}\})$. Since $\mathsf{OPT}_{1}^{\mathsf{B1}}$ exists but the condition in Algorithm 2 did not apply, we conclude that $\mathsf{opt}_{1}^{\mathsf{B1}}\geq\mathsf{opt}_{1}^{\min}+2=\mathsf{opt}_{1}^{\mathsf{C2}}+2$.
    Subtype $\mathsf{C2(i)}$:
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+4$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vy,vw\}$ and at most one more edge, which is guaranteed by Lemma 24 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, continuing (1) gives
    $\mathsf{red}(G)$ $\leq\mathsf{opt}_{1}-2+\alpha\cdot(\mathsf{opt}_{2}+4)+4\varepsilon\lvert V(G_{2}^{\mathsf{C2}})\rvert-4-2$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(4-\mathsf{opt}_{1})\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
    Subtype $\mathsf{C2(ii)}$:
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+5$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vz,zy,wy\}$ and at most one more edge, which is guaranteed by Lemma 24 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, (2) gives
    $\mathsf{red}(G)$ $\leq\mathsf{opt}_{1}-2+\alpha\cdot(\mathsf{opt}_{2}+5)+4\varepsilon\lvert V(G_{2}^{\mathsf{C2}})\rvert-4-3$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(5-\mathsf{opt}_{1})\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
    Subtype $\mathsf{C2(iii)}$:
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+4$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vy,wy\}$ and at most one more edge, which is guaranteed by Lemma 24 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, (3) gives
    $\mathsf{red}(G)$ $\leq\mathsf{opt}_{1}-2+\alpha\cdot(\mathsf{opt}_{2}+4)+4\varepsilon\lvert V(G_{2}^{\mathsf{C2}})\rvert-4-2$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(4-\mathsf{opt}_{1})\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
  3. (c)
    $(\mathsf{B2},\{\mathsf{A},\mathsf{B1},\mathsf{B2}\})$. In this case $\mathsf{OPT}_{1}^{\mathsf{B2}}$ exists but $t^{\min}=\mathsf{C2}$. Thus, the definition of Algorithm 2 and Definition 14 imply that $\mathsf{opt}_{1}^{\mathsf{C2}}\leq\mathsf{opt}^{\mathsf{B2}}_{1}-1=\mathsf{opt}_{1}-1$. Furthermore, in this case at least two vertices of $\{u,v,w\}$ are in a 2EC component in $\mathsf{OPT}_{2}$.
    Subtype $\mathsf{C2(i)}$:
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+3$ because we can turn $\mathsf{OPT}_{2}$ with two of the dummy edges $\{uy,vy,vw\}$ and at most one more edge, which is guaranteed by Lemma 24 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, (1) gives
    $\mathsf{red}(G)$ $\leq\mathsf{opt}_{1}-1+\alpha\cdot(\mathsf{opt}_{2}+3)+4\varepsilon\lvert V(G_{2}^{\mathsf{C2}})\rvert-4-2$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(3-\mathsf{opt}_{1})\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
    Subtype $\mathsf{C2(ii)}$:
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+4$ because we we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{vz,zy\}$, one of the dummy edges $\{uy,wy\}$, and at most one more edge, which is guaranteed by Lemma 24 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, (2) gives
    $\mathsf{red}(G)$ $\leq\mathsf{opt}_{1}-1+\alpha\cdot(\mathsf{opt}_{2}+4)+4\varepsilon\lvert V(G_{2}^{\mathsf{C2}})\rvert-4-3$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(4-\mathsf{opt}_{1})\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
    Subtype $\mathsf{C2(iii)}$:
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+3$ because we can turn $\mathsf{OPT}_{2}$ with two of the dummy edges $\{uy,vy,wy\}$ and at most one more edge, which is guaranteed by Lemma 24 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, (3) gives
    $\mathsf{red}(G)$ $\leq\mathsf{opt}_{1}-1+\alpha\cdot(\mathsf{opt}_{2}+3)+4\varepsilon\lvert V(G_{2}^{\mathsf{C2}})\rvert-4-2$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(3-\mathsf{opt}_{1})\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
  4. (d)
    $(\mathsf{C1},\{\mathsf{A},\mathsf{B1},\mathsf{C1}\})$. In this case $\mathsf{opt}_{1}^{\mathsf{C1}}$ exists but $t^{\min}=\mathsf{C2}$. Thus, the definition of Algorithm 2 and Definition 14 imply that $\mathsf{opt}_{1}^{\mathsf{C2}}\leq\mathsf{opt}^{\mathsf{C1}}_{1}-1=\mathsf{opt}_{1}-1$.
    Subtype $\mathsf{C2(i)}$:
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+3$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vy,vw\}$ into a solution for $G_{2}^{\mathsf{C2}}$. Thus, continuing (1) gives
    $\mathsf{red}(G)$ $\leq\mathsf{opt}_{1}-1+\alpha\cdot(\mathsf{opt}_{2}+3)+4\varepsilon\lvert V(G_{2}^{\mathsf{C2}})\rvert-4-2$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(3-\mathsf{opt}_{1})\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
    Subtype $\mathsf{C2(ii)}$:
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+4$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vz,zy,wy\}$ into a solution for $G_{2}^{\mathsf{C2}}$. Thus, continuing (2) gives
    $\mathsf{red}(G)$ $\leq\mathsf{opt}_{1}-1+\alpha\cdot(\mathsf{opt}_{2}+4)+4\varepsilon\lvert V(G_{2}^{\mathsf{C2}})\rvert-4-3$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(4-\mathsf{opt}_{1})\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
    Subtype $\mathsf{C2(iii)}$:
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+3$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vy,wy\}$ into a solution for $G_{2}^{\mathsf{C2}}$. Thus, continuing (3) gives
    $\mathsf{red}(G)$ $\leq\mathsf{opt}_{1}-1+\alpha\cdot(\mathsf{opt}_{2}+3)+4\varepsilon\lvert V(G_{2}^{\mathsf{C2}})\rvert-4-2$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(3-\mathsf{opt}_{1})\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
  5. (e)
    $(\mathsf{C2},\{\mathsf{A},\mathsf{B1}\})$. We have $\mathsf{opt}_{1}^{\mathsf{C2}}=\mathsf{opt}_{1}$.
    Subtype $\mathsf{C2(i)}$:
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+3$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vy,vw\}$ into a solution for $G_{2}^{\mathsf{C2}}$. Thus, continuing (1) gives
    $\mathsf{red}(G)$ $\leq\mathsf{opt}_{1}+\alpha\cdot(\mathsf{opt}_{2}+3)+4\varepsilon\lvert V(G_{2}^{\mathsf{C2}})\rvert-4-2$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(2-\mathsf{opt}_{1})+\alpha\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
    Subtype $\mathsf{C2(ii)}$:
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+4$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vz,zy,wy\}$ into a solution for $G_{2}^{\mathsf{C2}}$. Thus, continuing (2) gives
    $\mathsf{red}(G)$ $\leq\mathsf{opt}_{1}+\alpha\cdot(\mathsf{opt}_{2}+4)+4\varepsilon\lvert V(G_{2}^{\mathsf{C2}})\rvert-4-3$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(3-\mathsf{opt}_{1})+\alpha\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
    Subtype $\mathsf{C2(iii)}$:
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+3$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vy,wy\}$ into a solution for $G_{2}^{\mathsf{C2}}$. Thus, continuing (3) gives
    $\mathsf{red}(G)$ $\leq\mathsf{opt}_{1}+\alpha\cdot(\mathsf{opt}_{2}+3)+4\varepsilon\lvert V(G_{2}^{\mathsf{C2}})\rvert-4-2$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(2-\mathsf{opt}_{1})+\alpha\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
  6. (f)
    $(\mathsf{C3},\mathsf{A})$. We have $\mathsf{opt}_{1}^{\mathsf{C2}}=\mathsf{opt}_{1}^{\min}\leq\mathsf{opt}_{1}^{\mathsf{C3}}=\mathsf{opt}_{1}$.
    Subtype $\mathsf{C2(i)}$:
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+2$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vy\}$ into a solution for $G_{2}^{\mathsf{C2}}$. Thus, continuing (1) gives
    $\mathsf{red}(G)$ $\leq\mathsf{opt}_{1}+\alpha\cdot(\mathsf{opt}_{2}+2)+4\varepsilon\lvert V(G_{2}^{\mathsf{C2}})\rvert-4-2$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(2-\mathsf{opt}_{1})\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
    Subtype $\mathsf{C2(ii)}$:
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+3$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vz,zy\}$ into a solution for $G_{2}^{\mathsf{C2}}$. Thus, continuing (2) gives
    $\mathsf{red}(G)$ $\leq\mathsf{opt}_{1}+\alpha\cdot(\mathsf{opt}_{2}+3)+4\varepsilon\lvert V(G_{2}^{\mathsf{C2}})\rvert-4-3$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(3-\mathsf{opt}_{1})\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
    Subtype $\mathsf{C2(iii)}$:
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+2$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vy\}$ into a solution for $G_{2}^{\mathsf{C2}}$. Thus, continuing (3) gives
    $\mathsf{red}(G)$ $\leq\mathsf{opt}_{1}+\alpha\cdot(\mathsf{opt}_{2}+2)+4\varepsilon\lvert V(G_{2}^{\mathsf{C2}})\rvert-4-2$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(2-\mathsf{opt}_{1})\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
(Algorithm 2: $t^{\min}=\mathsf{C3}$)
Let $D^{\mathsf{C3}}\coloneq\{uv,uv,vw,vw\}$ denote the set of dummy edges used for the construction of $G_{2}^{\mathsf{C3}}$. First observe that the definition of Algorithm 2, Lemma 33, and the induction hypothesis gives
$\mathsf{red}(G)=|H^{\mathsf{C3}}|$ $\leq|\mathsf{OPT}_{1}^{\mathsf{C3}}|+\lvert H_{2}^{\mathsf{C3}}\rvert+\lvert F^{\mathsf{C3}}\rvert\leq\mathsf{opt}_{1}^{\mathsf{C3}}+\mathsf{red}(G_{2}^{\mathsf{C3}})$
$\leq\mathsf{opt}_{1}^{\mathsf{C3}}+\alpha\cdot\mathsf{opt}(G_{2}^{\mathsf{C3}})+4\varepsilon\lvert V(G_{2}^{\mathsf{C3}})\rvert-4\ .$ (4)
We distinguish the following cases depending on which type combination of Proposition 18 is present for $(\mathsf{OPT}_{1},\mathsf{OPT}_{2})$.
  1. (a)
    $(\mathsf{A},\{\mathsf{A},\mathsf{B1},\mathsf{B2},\mathsf{C1},\mathsf{C2},\mathsf{C3}\})$. Since $\mathsf{OPT}_{1}^{\mathsf{A}}$ exists, Lemma 22 gives $\mathsf{opt}_{1}=\mathsf{opt}_{1}^{\mathsf{A}}\geq\mathsf{opt}_{1}^{\mathsf{C3}}+3$. Moreover, $\mathsf{opt}(G_{2}^{\mathsf{C3}})\leq\mathsf{opt}_{2}+4$ because we can turn $\mathsf{OPT}_{2}$ with at most four dummy edges of $D^{\mathsf{C3}}$ into a solution for $G_{2}^{\mathsf{C3}}$. Thus, (4) gives
    $\mathsf{red}(G)$ $\leq\mathsf{opt}_{1}-3+\alpha\cdot(\mathsf{opt}_{2}+4)+4\varepsilon\lvert V(G_{2}^{\mathsf{C3}})\rvert-4$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(3-\mathsf{opt}_{1})+\alpha\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
  2. (b)
    $(\mathsf{B1},\{\mathsf{A},\mathsf{B1},\mathsf{B2},\mathsf{C1},\mathsf{C2}\})$. Since $\mathsf{OPT}_{1}^{\mathsf{B1}}$ exists but the condition in Algorithm 2 did not apply, we conclude that $\mathsf{opt}_{1}=\mathsf{opt}_{1}^{\mathsf{B1}}\geq\mathsf{opt}_{1}^{\min}+2=\mathsf{opt}_{1}^{\mathsf{C3}}+2$. Moreover, $\mathsf{opt}(G_{2}^{\mathsf{C3}})\leq\mathsf{opt}_{2}+3$ because we can turn $\mathsf{OPT}_{2}$ with at most three dummy edges of $D^{\mathsf{C3}}$ into a solution for $G_{2}^{\mathsf{C3}}$. Thus, (4) gives
    $\mathsf{red}(G)$ $\leq\mathsf{opt}_{1}-2+\alpha\cdot(\mathsf{opt}_{2}+3)+4\varepsilon\lvert V(G_{2}^{\mathsf{C3}})\rvert-4$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(2-\mathsf{opt}_{1})+\alpha\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
  3. (c)
    $(\mathsf{B2},\{\mathsf{A},\mathsf{B1},\mathsf{B2}\})$. In this case $\mathsf{OPT}_{1}^{\mathsf{B2}}$ exists but $t^{\min}=\mathsf{C3}$. Thus, the definition of Algorithm 2 and Definition 14 imply that $\mathsf{opt}_{1}^{\mathsf{C3}}\leq\mathsf{opt}^{\mathsf{B2}}_{1}-1=\mathsf{opt}_{1}-1$. Moreover, $\mathsf{opt}(G_{2}^{\mathsf{C3}})\leq\mathsf{opt}_{2}+2$ because we can turn $\mathsf{OPT}_{2}$ with at most two dummy edges of $D^{\mathsf{C3}}$ into a solution for $G_{2}^{\mathsf{C3}}$. Thus, (4) gives
    $\mathsf{red}(G)$ $\leq\mathsf{opt}_{1}-1+\alpha\cdot(\mathsf{opt}_{2}+2)+4\varepsilon\lvert V(G_{2}^{\mathsf{C3}})\rvert-4$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(1-\mathsf{opt}_{1})+\alpha\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
  4. (d)
    $(\mathsf{C1},\{\mathsf{A},\mathsf{B1},\mathsf{C1}\})$. In this case $\mathsf{OPT}_{1}^{\mathsf{C1}}$ exists but $t^{\min}=\mathsf{C3}$. Thus, the definition of Algorithm 2 and Definition 14 imply that $\mathsf{opt}_{1}^{\mathsf{C3}}\leq\mathsf{opt}^{\mathsf{C1}}_{1}-1=\mathsf{opt}_{1}-1$. Moreover, $\mathsf{opt}(G_{2}^{\mathsf{C3}})\leq\mathsf{opt}_{2}+2$ because we can turn $\mathsf{OPT}_{2}$ with at most two dummy edges of $D^{\mathsf{C3}}$ into a solution for $G_{2}^{\mathsf{C3}}$. Thus, (4) gives
    $\mathsf{red}(G)$ $\leq\mathsf{opt}_{1}-1+\alpha\cdot(\mathsf{opt}_{2}+2)+4\varepsilon\lvert V(G_{2}^{\mathsf{C3}})\rvert-4$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(1-\mathsf{opt}_{1})+\alpha\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
  5. (e)
    $(\mathsf{C2},\{\mathsf{A},\mathsf{B1}\})$. In this case $\mathsf{OPT}_{1}^{\mathsf{C2}}$ exists but $t^{\min}=\mathsf{C3}$. Thus, the definition of Algorithm 2 and Definition 14 imply that $\mathsf{opt}_{1}^{\mathsf{C3}}\leq\mathsf{opt}^{\mathsf{C2}}_{1}-1=\mathsf{opt}_{1}-1$. Moreover, $\mathsf{opt}(G_{2}^{\mathsf{C3}})\leq\mathsf{opt}_{2}+1$ because we can turn $\mathsf{OPT}_{2}$ with at most one dummy edge of $D^{\mathsf{C3}}$ into a solution for $G_{2}^{\mathsf{C3}}$. Thus, (4) gives
    $\mathsf{red}(G)$ $\leq\mathsf{opt}_{1}-1+\alpha\cdot(\mathsf{opt}_{2}+1)+4\varepsilon\lvert V(G_{2}^{\mathsf{C3}})\rvert-4$
    $\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4+(\alpha-1)(1-\mathsf{opt}_{1})\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
  6. (f)
    $(\mathsf{C3},\mathsf{A})$. Using $\mathsf{opt}_{1}^{\mathsf{C3}}=\mathsf{opt}_{1}$ and $\mathsf{opt}(G_{2}^{\mathsf{C3}})\leq\mathsf{opt}_{2}$ as $\mathsf{OPT}_{2}$ is already feasible for $G_{2}^{\mathsf{C3}}$, (4) implies
    $\mathsf{red}(G)\leq\mathsf{opt}_{1}+\alpha\cdot\mathsf{opt}_{2}+4\varepsilon\lvert V(G_{2}^{\mathsf{C3}})\rvert-4\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\lvert V(G)\rvert-4\ .$
This completes the proof of the lemma. ∎
Appendix B Canonical 2-Edge Cover
This section is dedicated to the proof of Definition 7, which we first restate.
[…]
Proof.
We first compute a triangle-free $2$-edge cover $H^{\prime}$ (which can be done in polynomial time due to Lemma 4) and show how to modify it to obtain a canonical $2$-edge cover with the same number of edges. Since $H^{\prime}$ is a triangle-free $2$-edge cover, each component is either a 2EC component that contains at least $4$ edges, or complex.
If $H^{\prime}$ is canonical, we are done. Otherwise, we apply the following algorithm to $H^{\prime}$: While there exist sets of edges $F_{A}\subseteq E\setminus H^{\prime}$ and $F_{R}\subseteq H^{\prime}$ with $|F_{A}|\leq|F_{R}|\leq 2$ such that $H^{\prime\prime}\coloneq(H^{\prime}\setminus F_{R})\cup F_{A}$ is a triangle-free $2$-edge cover and either $|H^{\prime\prime}|<|H^{\prime}|$, or $|H^{\prime\prime}|=|H^{\prime}|$ and $H^{\prime\prime}$ contains strictly less connected components than $H^{\prime}$, or $H^{\prime\prime}$ contains strictly less bridges than $H^{\prime}$, or $H^{\prime\prime}$ contains less cut vertices in 2EC components than $H^{\prime}$, we update $H^{\prime}\coloneq H^{\prime\prime}$.
Clearly, each iteration of this algorithm can be executed in polynomial time. Moreover, there can only be polynomially many iterations. Thus, let $H$ be equal to $H^{\prime}$ when the algorithm terminates. We prove that $H$ is a canonical 2-edge cover of $G$. This proves the lemma because the algorithm does not increase the number of edge compared to a minimum triangle-free $2$-edge cover of $G$. To this end, we consider the following cases.
Case 1: $H$ contains a 2EC component $C$ such that $4\leq|E(C)|\leq 7$ and $C$ is not a cycle on $|E(C)|$ vertices. We distinguish three exhaustive cases, and prove in any case that Case 1 cannot occur.
Case 1.1: $C$ has a cut vertex $v$. Observe that $v$ must have degree $4$ in $C$. Furthermore, each connected component in $C\setminus\{v\}$ contains at least $2$ vertices, as otherwise $C$ contains parallel edges, a contradiction to $G$ being structured. There must be a connected component $K_{1}$ in $C\setminus\{v\}$ such that $K_{1}$ contains at most 2 vertices (and hence at most 1 edge). This is true as otherwise each connected component has at least 2 edges in $C\setminus\{v\}$ and since there are at least $2$ connected components in $C\setminus\{v\}$, because $v$ is a cut vertex, $C$ must contain at least $8$ edges, a contradiction. Assume that there is a component $K_{1}$ with $2$ vertices and let $u_{1}$ and $u_{2}$ be these vertices. Note that the three edges $\{vu_{1},u_{1}u_{2},u_{2}v\}$ must be in $C$ as otherwise $G$ is not structured since it must contain parallel edges. If there exists an edge $u_{1}w\in E\setminus H$ (or $u_{2}w\in E$ by symmetry) for $w\in V\setminus V(C)$, then we can remove $vu_{1}$ from $H$ and add $u_{1}w$ to obtain a triangle-free $2$-edge cover $H^{\prime\prime}$ with strictly less components and $|H|=|H^{\prime\prime}|$, a contradiction to the algorithm. If there exists an edge $u_{1}w\in E\setminus H$ (or $u_{2}w\in E$ by symmetry) for $w\in V(C)\setminus\{v,u_{1},u_{2}\}$, then we can remove $vu_{1}$ from $H$ and add $u_{1}w$ to obtain a triangle-free $2$-edge cover $H^{\prime\prime}$ with $|H^{\prime\prime}|=|H|$ in which $v$ is not a cut vertex anymore. If neither of these two cases apply, we conclude $\deg_{G}(u_{i})=2$ for $i=1,2$. Thus, $v-u_{1}-u_{2}-v$ is a contractible cycle, a contradiction to $G$ being structured.
Case 1.2: $C$ contains no 2-vertex cut (and, thus, $C$ is 3-vertex-connected ). This implies that every vertex has degree at least 3 in $C$, as otherwise the neighboring vertices of a vertex of $C$ with degree $2$ form a $2$-vertex cut. But then it must hold that $|V(C)|=4$, since $|V(C)|\leq 3$ is not possible with the above degree constraint and if $|V(C)|\geq 5$, then the number of edges in $C$ is at least $8$, a contradiction. But if $|V(C)|=4$ and $C$ is $3$-vertex-connected, $C$ must be a complete graph on $4$ vertices. Therefore, there is an edge $e\in E(C)$ such that $H\setminus\{e\}$ is a triangle-free 2-edge cover, a contradiction to the algorithm.
Case 1.3: $C$ contains a 2-vertex cut $\{v_{1},v_{2}\}$. Note that each connected component of $C\setminus\{v_{1},v_{2}\}$ must have an edge to both $v_{1}$ and $v_{2}$, as otherwise $C$ contains a cut vertex and we are in Case 1.1. Hence, there are at most $3$ connected components in $C\setminus\{v_{1},v_{2}\}$, as otherwise the number of edges in $C$ is at least 8.
First, assume that $C\setminus\{v_{1},v_{2}\}$ contains exactly $3$ connected components. Then there can be at most one connected component in $C\setminus\{v_{1},v_{2}\}$ containing 2 vertices, as otherwise the total number of edges in $C$ exceeds $7$. In any case, each vertex $u$ of some connected component in $C\setminus\{v_{1},v_{2}\}$ must have an edge $e_{u}$ to $\{v_{1},v_{2}\}$ as otherwise $C$ contains a cut vertex and we are actually in Case 1.1. If there is a vertex $u$ of some connected component in $C\setminus\{v_{1},v_{2}\}$ that has an edge $uw$ in $G$ to some vertex $w\in V\setminus V(C)$, then we can replace $e_{u}$ with $uw$ in $H$ and thereby decrease the number of connected components, a contradiction to the algorithm. Hence, assume that this is not true. Assume that there is exactly one connected component $K_{1}$ in $C\setminus\{v_{1},v_{2}\}$ containing 2 vertices $x_{1}$ and $x_{2}$. Note that $K_{1}$ must consist of the single edge $x_{1}x_{2}$ as otherwise we have at least $8$ edges in $C$. Let $K_{2}$ and $K_{3}$ be the other two connected components in $C\setminus\{v_{1},v_{2}\}$ consisting of single vertices $y$ and $z$, respectively. By the above arguments, we have that $C$ contains the edges $\{v_{1}x_{1},x_{1}x_{2},x_{2}v_{2},v_{1}y,yv_{2},v_{1}z,zv_{2}\}$ (up to symmetry). If there is an edge $x_{1}y$ in $G$ (or any of $\{x_{1}z,x_{2}y,x_{2}z,yz\}$ by symmetry), then we can remove $v_{2}y$ and $x_{1}v_{1}$ from $H$ and add $x_{1}y$ to obtain a triangle-free $2$-edge cover with strictly less edges than $H$, a contradiction to the algorithm. Hence, assume that none of the edges exists in $G$. But then all vertices in $\{x_{1},x_{2},y,z\}$ have degree $2$ in $G$ and, thus, $C$ is contractible, a contradiction. The case that there is no component in $C\setminus\{v_{1},v_{2}\}$ containing 2 vertices, i.e., in which each component consists of exactly a single vertex, is similar. This finishes the case that in $C\setminus\{v_{1},v_{2}\}$ there are exactly 3 connected components.
Hence, assume that $C\setminus\{v_{1},v_{2}\}$ contains exactly 2 connected components $K_{1}$ and $K_{2}$ and without loss of generality assume $|E(K_{1})|\geq|E(K_{2})|$. Recall that each connected component of $C\setminus\{v_{1},v_{2}\}$ must have an edge to both $v_{1}$ and $v_{2}$. Hence, $\{v_{1},v_{2}\}$ is incident to at least 4 edges in $C$. Thus, $|E(K_{2})|\leq|E(K_{1})|\leq 3$. If $|E(K_{1})|=3$, then $K_{1}$ is a path on 4 vertices. Otherwise, $K_{1}$ is a triangle on 3 vertices or a star on 4 vertices (one center and 3 leaves). If $K_{1}$ is a triangle, there must be an edge $e$ in $C$ such that $H\setminus\{e\}$ is a triangle-free $2$-edge cover, a contradiction to the algorithm. If $K_{1}$ is a star on 4 vertices, then there must be 3 edges from $K_{1}$ to $\{v_{1},v_{2}\}$, as otherwise $H$ is not a 2-edge cover. But then $C$ contains at least 8 edges, a contradiction. Hence, if $|E(K_{1})|=3$, then $K_{1}$ is a path on 4 vertices. But then one leaf of $K_{1}$ must be adjacent to $v_{1}$ in $C$ and the other leaf of $K_{1}$ must be adjacent to $v_{2}$ in $C$. Since $K_{2}$ is a single vertex in this case and must be incident to both $v_{1}$ and $v_{2}$, $C$ is a cycle on $7$ vertices, a contradiction. Hence, we have that $|E(K_{2})|\leq|E(K_{1})|\leq 2$. Since $G$ does not contain parallel edges, $K_{1}$ must be a simple path on $i$ vertices for some $i\in\{1,2,3\}$. The same is true for $K_{2}$. By similar reasoning as before we can observe that the leaves of $K_{1}$ and $K_{2}$, respectively, must be incident to $v_{1}$ and $v_{2}$ (or, if $K_{1}$ or $K_{2}$ is a single vertex, then it must be incident to both $v_{1}$ and $v_{2}$) and therefore we observe that $C$ is a simple cycle, a contradiction.
Case 2: There is some complex component $C$ of $H$ containing a pendant block $B$ with less than 6 edges. Assume there is a pendant block $B$ with exactly 5 edges. The cases in which $B$ has 3 or 4 edges is similar. Note that $B$ must contain either 4 or 5 vertices. If $B$ contains exactly $4$ vertices then $B$ must contain a simple cycle on 4 vertices with an additional edge $e$ being a chord of this cycle, which is redundant, and, thus, a contradiction to the algorithm. Hence, assume that $B$ contains exactly $5$ vertices, i.e., $B$ is a cycle $b_{1}-b_{2}-b_{3}-b_{4}-b_{5}-b_{1}$, and assume that $b_{1}$ is incident to the unique bridge $f$ in $C$. If either $b_{2}$ (or $b_{5}$ by symmetry) is incident to an edge $e$ with neighbor in $V\setminus V(B)$ we can remove $b_{1}b_{2}$ from $H$ and add $e$ to obtain a triangle-free $2$-edge cover $H^{\prime\prime}$. If $e$ is incident to some vertex in $V\setminus V(C)$, the number of connected components in $H^{\prime\prime}$ is strictly less compared to $H$, and otherwise, that is, $e$ is incident to some vertex in $V(C)\setminus V(B)$, or $H^{\prime\prime}$ has strictly less bridges compared to $H$. This is a contradiction to the algorithm. Thus, in the following we can assume that $b_{2}$ and $b_{5}$ are only adjacent to vertices of $V(B)$. Since $G$ is structured, $b_{3}$ and $b_{4}$ both must be incident to some edge $e_{3}$ and $e_{4}$ going to some vertex in $V\setminus V(B)$, respectively, as otherwise $B$ is contractible. Moreover, there must be the edge $b_{2}b_{5}$ in $G$, as otherwise $B$ is contractible. But then we can remove $b_{2}b_{3}$ and $b_{1}b_{5}$ from $H$ and add $b_{2}b_{5}$ and $e_{3}$ to obtain a triangle-free $2$-edge cover $H^{\prime\prime}$. If $e_{3}$ is incident to some vertex $V\setminus V(C)$, the number of connected components in $H^{\prime\prime}$ is strictly less compared to $H$, and otherwise, that is, $e_{3}$ is incident to some vertex $V(C)\setminus V(B)$, $H^{\prime\prime}$ has strictly less bridges compared to $H$. This is a contradiction to the algorithm.
Case 3: There is some complex component $C$ of $H$ containing a non-pendant block $B$ that contains less than $4$ edges, that is, $B$ is a triangle. In this case, we observe that in any case there must be an edge $e$ in $B$ such that $H\setminus\{e\}$ is a triangle-free $2$-edge cover, a contradiction to the algorithm. ∎
Appendix C Bridge Covering
In this section, we prove our main lemma for bridge covering, which we first restate.
[…]
We prove this lemma by exhaustively applying the following lemma, which is essentially proved in [18] in Appendix D.1.
Lemma 37 (Lemma 2.10 in [18]).
Let $H$ be a canonical $2$-edge cover of some structured graph $G$ containing at least one bridge. There is a polynomial-time algorithm that computes a canonical 2-edge cover $H^{\prime}$ of $G$ such that the number of bridges in $H^{\prime}$ is strictly less than the number of bridges in $H$ and $\mathrm{cost}(H^{\prime})\leq\mathrm{cost}(H)$.
For the proof of this lemma it is crucial that each connected component has a credit of at least 1. Note that in the work of Garg et al. [18], they might have components in their 2-edge cover, which are triangles. They need to assign credit such that each triangle also receives a credit of 1. However, in our case, this is not an issue as we do not have components that correspond to triangles.
Furthermore, in their proof, Garg et al. [18] shortly argue why, after one of these steps of the bridge covering, the resulting 2-edge cover remains canonical. In [18], the definition of canonical is slightly different to our definition of canonical. However, it is straightforward to check that after such a bridge covering step, the resulting 2-edge cover is also canonical with respect to our definition.